### v1

# shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation

Abstract
Unlike classical lexical overlap metrics such as BLEU, most current evaluation metrics for machine translation (for example, COMET or BERTScore) are based on black-box large language models. They often achieve strong correlations with human judgments, but recent research indicates that the lower-quality classical metrics remain dominant, one of the potential reasons being that their decision processes are more transparent. To foster more widespread acceptance of novel high-quality metrics, explainability thus becomes crucial. In this concept paper, we identify key properties as well as key goals of explainable machine translation metrics and provide a comprehensive synthesis of recent techniques, relating them to our established goals and properties. In this context, we also discuss the latest state-of-the-art approaches to explainable metrics based on generative models such as ChatGPT and GPT4. Finally, we contribute a vision of next-generation approaches, including natural language explanations. We hope that our work can help catalyze and guide future research on explainable evaluation metrics and, mediately, also contribute to better and more transparent machine translation systems.
1 Introduction
The field of evaluation metrics for Natural Language Generation (NLG), especially machine translation (MT) is in a crisis (Marie et al. 2021).111We published an earlier version of this paper (Leiter et al. 2022a) under a different title. Both versions consider the conceptualization of explainable metrics and are overall similar. However, while the old version contains several novel experiments, this new version puts a stronger emphasis on the survey of approaches for the explanation of MT metrics including the latest LLM based approaches. For example, this comprises techniques that return fine-grained error labels or natural language explanations. Despite the development of multiple high-quality evaluation metrics in recent years (Zhao et al. 2019; Zhang et al. 2020a; Rei et al. 2020; Sellam et al. 2020; Yuan et al. 2021), the Natural Language Processing (NLP) community appears hesitant to adopt them for assessing NLG systems (Marie et al. 2021; Gehrmann et al. 2023). Empirical investigations of Marie et al. (2021) indicate that the majority of MT papers relies on surface-level evaluation metrics such as BLEU and METEOR (Papineni et al. 2002; Banerjee and Lavie 2005), which were created two decades ago, a trend that may allegedly have worsened in recent times. These surface-level metrics cannot (even) measure semantic similarity of their inputs and are thus fundamentally flawed, particularly when it comes to assessing the quality of recent state-of-the-art MT systems (e.g. Peyrard 2019; Freitag et al. 2022), raising concerns about the credibility of the scientific field. We argue that the potential reasons for this neglect of recent high-quality metrics include: (i) non-enforcement by reviewers; (ii) easier comparison to previous research, for example, by copying BLEU-based results from tables of related work (potentially a pitfall in itself); (iii) computational inefficiency to run expensive new metrics at large scale; (iv) lack of trust in and transparency of high-quality black box metrics. In this work, we concern ourselves with the last-named reason, and address the explainability of such metrics.
In recent years, explainability has become a crucial research area in AI due to its potential benefits for users, designers, and developers of AI systems Samek et al. (2018); Vaughan and Wallach (2021).222 As of now, there is no universally accepted definition of explainability in the AI community. In this work, we adapt the definition by Barredo Arrieta et al. (2020), which we discuss in §2.2.
For users of the AI systems, explanations help them make more informed decisions (especially in high-stake domains) Sachan et al. (2020); Lertvittayakumjorn et al. (2021), better understand and hence gain trust of the AI systems Pu and Chen (2006); Toreini et al. (2020), and even learn from the AI systems to accomplish tasks more successfully Mac Aodha et al. (2018); Lai et al. (2020). For AI system designers and developers, explanations allow them to identify the problems and weaknesses of the systems Krause et al. (2016); Han et al. (2020), calibrate the confidence of the systems Zhang et al. (2020b), and improve them accordingly Kulesza et al. (2015); Lertvittayakumjorn and Toni (2021).
Explainability is particularly desirable for evaluation metrics. Sai et al. (2022) suggest that explainable NLG metrics should focus on providing more information than just a single score (such as fluency or adequacy). Celikyilmaz et al. (2020) stress the need for explainable evaluation metrics to spot system quality issues and to achieve a higher trust in the evaluation of NLG systems. Explanations indeed play a vital role in building trust for new evaluation metrics.333As an illustrating example, Moosavi et al. (2021) distrust using the metric BERTScore Zhang et al. (2020a) applied to a novel text generation task, as it assigns a score of 0.74 to a nonsensical output, which could be taken as an unreasonably high value. While BERTScore may indeed be unsuitable for their novel task, the score of 0.74 is meaningless here, as evaluation metrics may have arbitrary ranges. In fact, BERTScore typically has a particularly narrow range, so a score of 0.74 even for bad outputs may not be surprising (BERTScore provides a scaling method that solves these issues to some degree). Explainability techniques would be very helpful in preventing such misunderstandings, e.g. by showing that the features BERTScore attends to align with human judgement. For instance, one potential reason that a metric will likely be better accepted by the research community is if the explanations of the scores align well with human reasoning and faithfully reflect its internal decision process (see §2.2).
By contrast, if faithfulness is given and the explanations are counter-intuitive, users and developers might lower their trust and be alerted to take additional actions, such as trying to improve the metrics using insights from the explanations or looking for alternative metrics that are more trustworthy. Furthermore, explainable metrics can be used for other purposes: for example, when a metric produces a low score for a given input, highlighted words (a widely used method for explanation, see §4) in the input are natural candidates for manual post-editing.
This concept paper aims at providing a systematic overview of the existing efforts in explainable MT evaluation metrics and an outlook for promising future research directions. The main focus of this work lies on explainable evaluation metrics for MT, many of our observations and discussions can likely be adapted to other NLG tasks, however.
We make the following contributions, outlined by the structure of our paper:
§2 Background & Terminology: We give an overview of concepts important for explainable MT evaluation. We also highlight different goals and the audiences that require them.
§4 Taxonomy: We provide a structured overview of previous efforts in explainable MT evaluation. An overview is shown in Figure 4. The process of selecting works for this taxonomy is described in §3 Literature Review & Selection.
§5 Future Work: An analysis of underexplored research directions , and a collection of exemplary methods from other NLP domains to illustrate potential future paths. Here, we also discuss the usage of large language models (LLMs), like ChatGPT OpenAI (2023b) and GPT4 OpenAI (2023a).
With this work, we aim to solidify the field of explainable MT metrics and provide guidance for researchers, in specific metric developers, who aim to better understand and explain MT metrics, as well as users that want to employ explainability techniques to explain metric outputs. In the long term, we envision that our work will aid the development of improved MT metrics and thereby help to improve the quality of machine translations. Potentially, explanations can also aid other use cases such as translation selection and semi-automatic labeling (see §2.3). We also hope that our work will more generally inspire explainable NLG metric design.
2 Background & Terminology
In this section, we first introduce and relate definitions and dimensions of MT metrics and explainability, which we use in the later parts of the paper. Further, we collect goals and target audiences of explainable MT metrics.
2.1 Machine Translation Evaluation Metrics
MT metrics grade machine translated content, the hypothesis, based on ground truth data. We differentiate these metrics along several dimensions. A summary is shown in Table 1. We note that metrics can be categorized along many dimensions (e.g. Sai et al. 2022; Celikyilmaz et al. 2020). Here, we focus on a minimal specification that we will leverage later. Some parts of this section refer to explainability (see definitions in §2.2).
| Dimension | Description |
| Input type |
Whether source, reference translation or both are used as ground truth for comparison |
| Granularity |
At which level a metric operates: Word-level, sentence-level, document-level |
| Quality aspect |
What a metric measures: Adequacy, fluency, etc. |
| Learning objective |
How a metric is induced: Regression, ranking, etc. |
Input type We call a metric reference-based if it requires one or multiple human reference translations to compare with the hypothesis as ground truth, which can be seen as a form of supervision (e.g. Papineni et al. 2002; Zhao et al. 2019; Rei et al. 2022a). Reference-free metrics do not require a reference translation to grade the hypothesis (e.g. Zhao et al. 2020; Ranasinghe et al. 2020b; Belouadi and Eger 2023). Instead, they directly compare the source to the hypothesis. In the literature, reference-free MT evaluation is sometimes also referred to as “reference-less” (e.g. Mathur et al. 2020) or “quality estimation” (e.g. Zerva et al. 2022). This dimension is important for explainable MT evaluation, as explainability techniques might need to consider which resources are available.
Granularity Translation quality can be evaluated at different levels of granularity: word-level, sentence-level and document-level. The majority of metrics for MT return a single sentence-level score for each input sentence (e.g. Zhao et al. 2019; Zhang et al. 2020a). Beyond individual sentences, a metric may also score whole documents (multiple sentences) (Jiang et al. 2022; Zhao et al. 2023). In the MT community, metrics that evaluate translations at the word-level (for example, whether individual words are correct or not) are also common (Turchi et al. 2014; Shenoy et al. 2021). Recently, there has been a tendency to compare metrics to human scores derived from fine-grained error annotations — multidimensional quality metrics (MQM) (Lommel et al. 2014) — as they tend to correspond better to human professional translators (Freitag et al. 2021a). Metrics of higher granularity (e.g. word-level) provide more explanatory value than those of lower granularity, as they provide more details on which parts of the input might be translated incorrectly (e.g. Leiter 2021; Fomicheva et al. 2021, 2022). Most techniques we describe in §4 target the explainability of sentence-level metrics.
Quality aspect This refers to the properties a metric evaluates and measures, for example, human assigned adequacy (via direct assessment (Graham et al. 2016) obtained from crowd-workers), fluency, or other aspects (such as relevance, informativeness, or correctness, mostly in other NLG fields) (e.g. Sai et al. 2022; Celikyilmaz et al. 2020). We use this dimension in §4, as metrics that report multiple quality aspects provide more information on specific translations, i.e., they have more explanatory value for potential users.
Learning objective Yuan et al. (2021) describe a further distinction based on the task a metric is designed to solve (or how a metric is implemented) to achieve high correlation with human judgments. They identify the three tasks below.444Chen and Eger (2023) argue for metrics induced from natural language inference (NLI) as a conceptual framework.
Supervised Regression/Ranking: These models use supervised regression to train a metric to predict continuous values that directly correlate with human judgments from hypothesis and ground truth tokens. Alternatively, a metric may be trained to rank hypothesis sentences. For example, Rei et al. (2020) propose a ranking setup in which they minimize a loss that becomes smaller the greater the distance between embedding representations of the two sentences become. Especially in the MT community, metrics trained on human scores (e.g. COMET; Rei et al. 2022a) have become more and more dominant, while Belouadi and Eger (2023) argue that unsupervised metrics are more widely applicable.
Text Generation: Some metrics use the probability that a sentence (e.g., a translation) is created by a text generation system, given an input sentence (e.g., a reference sentence), as sentence-level score. Yuan et al. (2021) place the metric PRISM (Thompson and Post 2020) and their own metric BARTScore in this setup.
The distinction of learning objectives is important for explainable MT evaluation, as the applicability of some explainers depends on the objective.
MT metrics also have multiple use cases; see Figure 1 (in the following, preceding numbers correspond to this figure). (1) One use case is to rate a single translated text, for example, in a scenario where a human user wants to check their own human555Note, however, that humans might make different types of translation errors (Specia et al. 2018). Therefore, metric performance first needs to be evaluated for this use case. translation or a translation they received by an MT model before using it in a downstream task (e.g. Murgolo et al. 2022; Zouhar et al. 2023). (2) The standard use case is to rate a large number of translations to compare MT systems (e.g. Freitag et al. 2022). (3) Another use case is to employ metrics in the training process of MT models, e.g. using reinforcement learning (e.g. Wu et al. 2018) or optimizing by computing derivatives with Gumbel-softmax (e.g. Jauregi Unanue et al. 2021). (4) The fourth use case is to use the metrics in the decoding process of machine translation systems for reranking or with minimum Bayes risk decoding Fernandes et al. (2022a).
The use cases can have different audiences and explainability might play a different role to them (see 2.2). For example, anyone might be a user that wants to rate a translation, but usually only domain experts want to rate a translation model.
2.2 Explainability
In this section, we (1) define explainability in the scope of this work, (2) discuss the definitions and roles of faithfulness and plausibility for explainable MT metrics and (3) contrast explainability with the terms robustness and fairness. Then, we (4) discuss the properties we use to structure and describe the taxonomy of prior works on explainable MT metrics (see §4), namely explanation level, explanation type, model access and audiences. Based on (1), we decide on the scope of papers we include in the taxonomy in §3.
Explainability & interpretability Prior work has noted that the terms interpretability and explainability vary in definition and are sometimes used ambiguously throughout the literature (Barredo Arrieta et al. 2020; Jacovi and Goldberg 2020; Vilone and Longo 2021). In this paper, we follow the notion of passive interpretability and active explainability (for example, Barredo Arrieta et al. 2020). In particular, we adapt the definition of Barredo Arrieta et al. (2020) as follows:
An MT metric is explainable if itself or some external instance, an explainer, can actively provide explanations that make its decision process or its output clearer to a certain audience.
This definition varies from Barredo Arrieta et al. (2020) in two aspects: (1) It includes external explainers (also called post-hoc methods) and (2) it includes explanations of the metric’s output. We note that (2) is different from explaining a metric, as output explanations may not provide insights into how a model works. Still, we include (2) in our explainability definition for two reasons. One is that there have been two shared tasks and other existing works on metric explainability that consider the plausibility aspect of explanations only (see the following paragraph on faithfulness and plausibility) (e.g. Fomicheva et al. 2020, 2021; Zerva et al. 2022; Xu et al. 2023). Thereby, explanations these works consider might be explanations of the output that do not actually describe model internals. We want to accommodate these works that do not check faithfulness in our definition. The other reason is that we see use cases of output explanations of strong metrics, i.e., plausible-only explanations (see the following paragraph on faithfulness and plausibility).
In contrast, a metric is interpretable if it can be passively understood by humans (for example, decision trees and k-nearest neighbors) (Barredo Arrieta et al. 2020). Modern MT metrics are mostly based on embeddings and black box language models, making them non-interpretable. The papers we categorize in this work generate explanations for MT metrics and their outputs, i.e. they consider explainability.
[set/.style = text opacity = 1]
[…]
Faithfulness & plausibility An explanation is faithful if it accurately represents the reasoning process underlying a model’s prediction(s) (Jacovi and Goldberg 2020). In contrast, an explanation is plausible if it convinces its audience that a model’s prediction is correct (Jacovi and Goldberg 2021), i.e., if it aligns with human judgment. Both properties are not necessarily dependent on each other (Jacovi and Goldberg 2020). For example, an explanation of an MT metric score could make sense to a human without reflecting the metrics internal details and vice-versa. A degree of faithfulness is indispensable to all explainability works that actually want to reason about a model’s decision processes. Plausibility, on the other hand, might convince explainees of decision processes that do not reflect the model’s internals correctly, or incorrect model predictions. As such, optimizing for plausible explanations might be unethical and lead to deceiving outcomes (Jacovi and Goldberg 2021; Jin et al. 2023). However, in our taxonomy, we find that most current works on explainable MT metrics do not consider faithfulness and only evaluate plausibility (see §4), often leaving unknown whether and to which degree the model’s internal process is accurately described by explanations. While this is a decisive current shortcoming, we argue that explanations that may be only plausible can still benefit certain audiences and use cases of MT metrics that do not target understanding the metric: (1) We assume that current high-quality metrics already work well to judge translation quality in many settings (Freitag et al. 2022; Zerva et al. 2022). Explainers that are proven to have a high plausibility to human experts, as for example measured by Fomicheva et al. (2021); Zerva et al. (2022), will likely provide strong reasoning about the translation quality when provided with a correct score by such metrics. We caution that this process will have two points of failure: metric and explainer, hence the overall rate of failures will be higher than for each of them individually. Still, we predict that this process will work reasonably well to foster the goals of metric accessibility (see §2.3), as long as users are aware of possible failures (see future work §5), similar to the problem of hallucinations in LLMs (Ji et al. 2023b). (2) Language experts should be able to discern incorrect explanations. For them, explanations of quality scores (produced by high quality metrics) that strive to be plausible could provide means to faster evaluating and (if necessary) correcting translations, as encompassed by the goals of metric accessibility and semi-automatic labeling (see §2.3). (3) LLM’s have been shown to provide better results when being prompted to explain them (Wei et al. 2022). Here, more plausible explanations might lead to better scores and vice-versa (see §5). (4) Similar to (3), language models sometimes benefit from multitask learning (Radford et al. 2019), so employing automatically generated explanations might assist training better metrics. Note that (1) and (2) correspond to the first metric use case in Figure 1. To summarize, plausible but not faithful explanations could still be useful either to human users or intrinsically to a metric itself in several scenarios.
Fairness & robustness In Figure 2, we show the relationship of our definition of explainability to robustness and fairness. Robustness measures how well a model (here, a metric) can address execution errors, erroneous and malicious inputs and unseen data (e.g Li et al. 2023). In contrast, Li et al. (2023) define fairness as the goal that a model should not have biases that result in unfair treatment for certain groups of people. As umbrella term for these concepts, we use responsible AI Barredo Arrieta et al. (2020).666There are many related and partly conflicting definitions and terms in explainable AI research (e.g. Vilone and Longo 2021) and the given definitions are not yet firmly established in the literature. For example, Sharma et al. (2020) define explainability, robustness and fairness as exclusive properties. The main focus of our work is to build a taxonomy of methods for explainable MT metrics and the chosen definitions allow us to reasonably relate existing work. Barredo Arrieta et al. (2020) define responsible AI as a concept that encourages the systematic adoption of several AI principles, while also setting a focus on accountability and privacy. Methods that test for robustness and fairness can all be grouped into the given explainability definition, as their fulfillment requires explanations that will also help certain audiences to understand a model’s behavior better Barredo Arrieta et al. (2020).777Barredo Arrieta et al. (2020) group robustness under the term confidence. In the MT metrics literature, this overlap in terminology is apparent for techniques that check how robust metrics are to input perturbations. Here, some works frame these as explainability techniques (e.g. Sai et al. 2021; Karpinska et al. 2022), while others frame them as robustness tests (e.g. He et al. 2022; Chen and Eger 2023); see §4. In Figure 2, we only let robustness and fairness half overlap with explainability to highlight that the explanatory value increases with fine-grained explanations. E.g., robustness could be measured coarse-grained with a single score, such as the minimal number of adversarial perturbations required to successfully attack a metric, or fine-grained, by using predefined perturbation templates (e.g. Sai et al. 2021) or attackers that will attack with perturbations of specific properties (e.g. Chen and Eger 2023). While a single score can also explain whether metric behavior is robust, susceptibility scores to various properties allow to assess metrics in greater detail. The same is true for fairness (Sun et al. 2022). Another reason that we do not fully include robustness and fairness in the explainability definition is that explainability results from testing them, but vice-versa robustness and fairness are usually not given just by having an explainable model.
We note that common explanation types in the evaluation of robustness and fairness are counterfactual explanations, i.e., minimal perturbed input samples that change the output of a model to demonstrate its behavior Wachter et al. (2017); Freiesleben (2021). For fairness evaluation, counterfactuals from domains that carry a risk of being biased are used Sharma et al. (2020); Sun et al. (2022). For robustness, existing works check whether metric scores change as expected, when presented with counterfactuals in contrast to the original hypotheses (e.g. Sai et al. 2021; Chen and Eger 2023); see §4 for more details. Some works frame these counterfactuals as adversarial attacks (e.g. Sharma et al. 2020; Chen and Eger 2023), others as diagnostic tests(ets) (e.g. Karpinska et al. 2022) or challenge sets (e.g. Freitag et al. 2022).
Explanation level
Explainability
techniques can be distinguished into decision-
understanding and model-understanding (e.g. Gehrmann et al. 2020; Holzinger et al. 2022).
Decision-understanding methods are described as
explaining the decision process for a specific model output with respect to a specific input while model-understanding methods are described as
explaining general properties of a model that steer its decisions. Often, decision-understanding is also referred to as local and model-understanding as global (e.g. Lipton 2016; Guidotti et al. 2018; Doshi-Velez and Kim 2017; Danilevsky et al. 2020).
Explanation type Further, we distinguish explanation types, i.e., the output format of the explanations (e.g. Wiegreffe and Marasovic 2021). One example are natural language explanations that describe the model output or its internal decision process with free text. In the case of MT metrics, this could be a sentence like “The metric returned a score of 0.6 for the translation, as it correctly captures all details of the source sentence but has severe grammatical errors at positions […]”. There are many possible explanation types (e.g. Barredo Arrieta et al. 2020; Wiegreffe and Pinter 2019), such as, feature importance explanations or counterfactuals. For our taxonomy, we select those that occur in our target papers (see §3). The definition of each explanation type is paired with the description of methods in §4.
Model Access Most taxonomies of explainability techniques consider the required level of model access (e.g. Barredo Arrieta et al. 2020; Linardatos et al. 2021; Madsen et al. 2022). Here, techniques that require no access (that can work with any black box model) are called model-agnostic. Techniques that require access to model parameters are called model-specific, as only specific model architectures satisfy the requirements of the method. Due to the prevalent role of transformer architectures, many modern evaluation metrics use embeddings and attention mechanisms and allow for gradient computations. This allows to apply many techniques that are specific to these architectures to such metrics. A special type of model-specific explanations are those that are given by a model (here a metric) itself. We call these metrics “explainable by design”, as they are designed in a way that provides more details on their explanations or decision process. Examples for MT metrics in this category are unsupervised matching metrics like BERTScore Zhang et al. (2020a) and MoverScore Zhao et al. (2019), as they compute token-level similarity scores that can be used as explanations of the sentence-level score Leiter (2021); Fomicheva et al. (2021). Finally, there are also interpretable models which are easy to understand for humans such as Naive Bayes and decision trees Freitas (2014); Arya et al. (2019). We can normally obtain decision-understanding explanations of these models at the same time as predictions (for example, the corresponding path in the decision tree), while the models can also be considered explanations of themselves (for example, the trained decision tree itself). We do not use the dimension of interpretable models in our taxonomy (see §4) as none of the state-of-the-art metrics falls into this category. Model-agnostic techniques require least model access, model-specific techniques, which use the gradient, embeddings or attention weights require more access and models that are explainable by design require access to the full model to generate explanations.888Note that besides interpretable models, none of the access levels inherently leads to faithful explanations, i.e. explanations that correctly describe the internal processes of a model. Hence, this property needs to be considered separately (e.g. Jacovi and Goldberg 2020).
Audiences Prior works in explainability point out that different audiences require different types of explanations (e.g. Barredo Arrieta et al. 2020; Sharma et al. 2020). For example, non-experts might have difficulties understanding explanations given in the domain language of experts. Barredo Arrieta et al. (2020) distinguish model users, affected users, developers, regulatory entities and managers. Users of MT metrics are usually domain experts in MT that want to evaluate or improve MT systems, i.e. MT developers. Another group of users might be non-experts that want to use the models to check their own translations or machine generated translations. Affected users might be the ones that receive translations that were graded by a metric or generated by a model that was supervised or tested with the metric. For example, Fernandes et al. (2022a) employ MT metrics in the decoding process of MT models. We consider users of those MT models as affected users. Developers are the designers of MT metrics. Regulatory entities and managers have an interest in enforcing guidelines. We describe the relation of the audiences towards goals of explainable MT metrics in §2.3 and associate audiences with explanation types in §4.
2.3 Goals of Explainable MT Metrics
In this section, we discuss goals that require specific focus for explainable MT evaluation. We orient ourselves on common goals proposed by Lipton (2016); Barredo Arrieta et al. (2020); Jacovi and Goldberg (2021) and identify four specific ones that match the literature we identify in §3 and use cases that might commonly occur in our field of research (see Figure 3). We follow with a short description of each of them below.
Goal 1: Diagnose and improve metrics. By explaining why a metric predicted a certain score for a machine translation, where humans would assign a different score, developers might understand its weaknesses (e.g. Kaster et al. 2021; Sai et al. 2021). This can enable architectural changes or the selection of different data sets, leading to metric improvement. Likewise, explaining the whole metric can unveil whether it follows desired general properties or might otherwise be led astray by carefully crafted adversarial inputs. Goal 1 is especially important for the audience metric developers, as they need to understand the shortcomings to improve on them.
Goal 2: Make metrics more expressive and accessible. A metric that assigns a score on a sentence level is difficult to understand for non-experts, e.g., non-native speakers. If a metric provides further information, such as which words it considers incorrect, it may become more accessible and thus foster trust (which may lead to more widespread adoption of the metric). Exemplary works in this direction are Fomicheva et al. (2021); Bao et al. (2023). Accessible metrics might also help with use cases like translation error correction in the downstream tasks of language learning or human translation services (e.g. Zouhar et al. 2023). In these latter two use cases, accessibility might not require faithfulness of the explanation (see §2.2). Goal 2 is important for the audience of metric users. These can be either non-experts that want to understand specific metric outputs or developers of MT systems that want to understand if a metric is a good fit for their tasks (see Figure 1).
Goal 3: Support Semi-Automatic labeling. Fine-grained annotations like word-level translation error labels (e.g. Fomicheva et al. 2021) are especially difficult to obtain. Obtaining automatic explanations to aid human annotators could boost their efficiency tremendously (e.g. Desmond et al. 2021). To our best knowledge, none of the current works has applied explainers in these settings for metrics. Semi-automatic labeling could also help with post-editing of translations (e.g. Specia and Farzindar 2010) to guide how sentences should be adapted. Goal 3 is important for metric developers that want to create such word-level translation error data sets and metric users that want to correct translations.
Goal 4: Checking for social biases. Learned metrics might be biased towards certain types of texts used during training (e.g. Sun et al. 2022). Biases could be detected by observing explanations where sensitive attributes (for example, gender, political or racial aspects) are flagged to be important (Lipton 2016). For example, explanations could show that a metric considers male names more important for a translation than female names in the same scenario. Goal 4 can be important for all audiences. Affected users might be treated unfairly if biases occur in a metric. Metric users, metric developers, regulatory entities and managers might want to prevent this from happening to conform with given regulations.
3 Literature Review & Selection
To build the taxonomy in §4, we survey literature which either explicitly state that they address the explainability of MT metrics or propose techniques that make MT metrics or their outputs clearer to understand by certain audiences (techniques that provide explanations). We restrict our search and taxonomy to works that consider neural network based MT metrics. We conducted an initial literature search by selecting 11 “seed papers”. Then, we leveraged the API of Semantic Scholar999https://www.semanticscholar.org/ to query new papers citing the work of the seed papers and papers citing those papers (citation tree at depth 2). This amounted to 383 papers. Next, we filtered the papers by keywords, leaving 177 papers. Then, we manually annotated whether their contents fit our definitions. Thereby, we identified 12 papers that consider the explainability of MT metrics. In later stages, we progressively added further work released after the initial search and related works that we identified from other sources. In total, we describe 29 papers, some of which are not shown in Figure 4 due to space constraints.
Following §2.1, we consider higher granularity scores (e.g. word-level) as explanations for lower granularity metrics (e.g. sentence-level). However, without requiring faithfulness this definition allows to view any word-level metric as explainers for any sentence-level metric. Thereby, it would even include cases in which the word-level scores are determined completely independent from the sentence-level, i.e., cases where the explainer would not be faithful. Due to the high number of existing word-level approaches, for this explanation type, we only consider works that explicitly employ the word-level scores for explainability.
Likewise, following §2.1, metrics that report multiple quality aspects such as adequacy or fluency provide more explanatory value on why an explanation is good or bad than metrics that return a single score. For the quality aspect explanation type, we make an exemplary selection of techniques derived from recent metrics that leverage generation probabilities (e.g. Thompson and Post 2020; Yuan et al. 2021).
We also include a few non-MT works into our taxonomy, which show how some aspects that are currently not addressed by the MT related works might be tackled.
[…]
[category] (-0.100000-5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodequality
aspects (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodefeature
importance (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle nodefine-grained
errors (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000);
[category] (-0.100000-5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodeperturbation
robustness (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000) rectangle nodelinguistic
properties (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
[category] (0.100000- 5.100000/2, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node post-hoc (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeblack-box (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[title] (0.100000+ 5.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodegradient (0.100000+ 5.100000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeattention
weights (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeembeddings (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[category] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000- 5.100000/2, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node intrinsic (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000);
[…]
[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodemodel specific (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Thompson and Post 2020)
(Yuan et al. 2021)
(Kamal Eddine et al. 2023)
(Fu et al. 2023) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Kaster et al. 2021) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Golovneva et al. 2023)
(Opitz and Frank 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Rei et al. 2022b)
(Rei et al. 2023) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000) rectangle node(Treviso et al. 2021) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Fernandes et al. 2022b)
(Kabir and Carpuat 2021)
(Rubino et al. 2021)
(Wang et al. 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Leiter 2021)
(Tao et al. 2022)
(Azadi et al. 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Eksi et al. 2021) (0.100000+ 5.100000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Fomicheva et al. 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Leiter et al. 2022b) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node(Lu et al. 2022, 2023)
(Xu et al. 2023), … (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Sai et al. 2022)
(Karpinska et al. 2022)
(He et al. 2022)
(Sun et al. 2022)
(Chen and Eger 2023)
… (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
[content] (0.100000, 0.100000+ 0.985000+ 0.100000) rectangle node(Kaster et al. 2021) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000) rectangle node(Opitz and Frank 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
[-] (0.100000- 0.100000- 5.100000/1.75 + 0.3, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) – (0.100000- 0.100000- 5.100000/1.75 - 0.100000+0.3, 0.100000+ 0.985000+ 0.100000) node[anchor=mid, below, rotate=-90, xshift=-6cm] decision-understanding model-understanding;
[set/.style = draw]
[…]
Adequacy: 0.7
Fluency: 1;
[…]
“Things became complicated”
“Die Dinge bekamen kompliziert”;
[…]
“Things sensebecame/sense complicated”
“Die Dinge sensebekamen/sense kompliziert”;
[…]
Numeric: “3 Dinge bekamen kompliziert”lower score
Order: “kompliziert die Dinge bekamen”lower score
[…]
(Example)[set, align=left, rectangle, dotted] at (-6.5, 5.8) Example Data:
Ground Truth: “Things became complicated”
Hypothesis: “Die Dinge bekamen kompliziert”
Score: ;
[…]
4 Taxonomy
Based on the background described in the previous sections and the literature search we conducted, we create a taxonomy of prior works that consider the explainability of MT metrics. This taxonomy is shown in Figure 4. The taxonomy contains five explanation types based on the selected papers’ contents: quality aspects, feature importance, fine-grained errors, perturbation robustness and linguistic properties. We structure this section based on these explanation types. Figure 5 shows an example for each explanation type. For each type, we describe the techniques applied by the previous works and how they relate to each other. We begin with explanation types that explain specific samples and move to techniques that explain metric behavior in general.
Quality aspects (Row 1 in Figure 4) This category comprises works that present explanations of the metric output by producing separate scores for different aspects of the translation. Most methods that we list develop generation-based metrics that are explainable by design (“model specific” in Figure 4). This means, they do not use an external explainer that could be applied to other metrics. BARTScore (Yuan et al. 2021) uses the autoregressive generation model BART (Lewis et al. 2020) to predict the average word generation probability of a sentence , given a sentence :
This formula sums over the log-probability of token being generated by BART with parameters given all previous tokens of and all tokens of . Depending on whether A is the ground truth sentence and B the hypothesis (or vice-versa), Yuan et al. (2021) state that different aspects of the translation are evaluated. For example, they associate the generation direction of A:source B:hypothesis with the properties coherence and fluency. This idea of using the generation probabilities in multiple directions was first introduced in PRISM (Thompson and Post 2020), but Yuan et al. (2021) formalized the connection of the directions to known quality aspects. DATScore (Kamal Eddine et al. 2023) extends BARTScore by using a multilingual model and further generation directions (additionally using the source sentence). The metric GPTScore (Fu et al. 2023) provides a suite of prompts for language models that emphasize specific aspects when requesting a translation or paraphrase. Then, similar to BARTScore, the generation probabilities are averaged for each word. For example, they use the prompt “Rewrite the following text to make it more grammatical and well-written: [reference or hypothesis] In other words,” to generate a paraphrase whose probabilities they use for a fluency score. The example specifies reference/hypothesis to describe that the metric checks either the generation probability of the hypothesis given the reference, or vice-versa.
Another way to gather fine-grained scores is to apply suites of metrics (each of which considers different aspects), like ROSCOE Golovneva et al. (2023), an embedding based metric suite, which has been proposed to grade step-by-step reasoning processes (column “embeddings” in Figure 4).
Besides adding explanatory value per se, quality aspect explanations have two more potential use cases for explainable metrics: (1) separate quality scores could be aggregated, to compute single, decomposable sentence-level scores and (2) scores could be aggregated over data sets to check which aspects a metric favors the most. For example, Kaster et al. (2021); Opitz and Frank (2022) also assign quality aspect scores, but mostly use them for model understanding; see linguistic properties explanations for more details.
In the example presented in Figure 5, we report fluency and adequacy as quality aspects. We assign a fluency score of 1 on a scale from 0 to 1, as the translation is fluent and an adequacy of 0.7 as the translation of “became” is incorrect.
Goals: Quality aspect approaches could be used for metric accessibility and semi-automatic labeling. The usage for metric diagnosis is limited, as the approaches are often based on models that are explainable by design and might not be faithful (the models generate their own explanations). The current approaches cannot be used to detect social biases.
Audiences: The quality aspects fluency and adequacy follow simple definitions, such that these metrics could be used by non-expert and expert users.
Limitations: The generation directions might not fully match the quality aspects. Also, the usage of prompting leaves it to another black box setting whether the model picks up the desired properties or not. Lastly, the explanatory value of quality aspect approaches per se is rather small, as they only allow to reason, “the rated translation is good because of aspect x” and to guess about the model’s reasoning. As none of the current works101010Kaster et al. (2021); Opitz and Frank (2022) are considered at a later point. specifically targets explainability, they do not consider faithfulness. Plausibility is automatically considered in standard evaluation schemes with human judgments.
Feature importance (Row 2 in Figure 4) Feature importance techniques assign scores to all features in the input of a metric (here the input tokens) that express their importance to the output of the metric.111111Other names include relevance scores or attribution scores. Sometimes the scores are used to highlight important words. Fomicheva et al. (2022) note that high feature importance scores should correspond to translation errors, as humans pay most attention to errors when rating a translation. They build on the idea that when humans evaluate translations, they often focus on the errors that they can identify on a word- or phrase-level (Freitag et al. 2021a). Then, Fomicheva et al. (2022) evaluate how well feature importance is correlated with human word-level error annotations, i.e., the plausibility of the explanations.
We note that, depending on the explainability method, these scores might be inverted, following the interpretation “errors are not important for a metric to achieve a high score”.121212Most gradient and attention based approaches return high scores for errors (e.g. Treviso et al. 2021; Rubino et al. 2021), for embedding-based approaches it depends on whether cosine-similarity or -distance is used (e.g. Leiter 2021; Tao et al. 2022), for attribution based techniques using LIME (Ribeiro et al. 2016) or SHAP (Lundberg and Lee 2017), low scores are returned for errors (e.g., as part of the baseline implementations in Eval4NLP21: https://github.com/eval4nlp/SharedTask2021/tree/main/baselines). Two shared tasks for the explainabiltiy of reference-free metrics have originated from the setting in Fomicheva et al. (2022) so far (Fomicheva et al. 2021; Zerva et al. 2022). As the respective papers already give an overview of the participating methods, we will only briefly describe the methods and set our focus to describe the model access they require. We describe the model access levels of Figure 4 from left to right. Approaches that work directly on black box models are explored in Fomicheva et al. (2022); Eksi et al. (2021); Treviso et al. (2021). They leverage model agnostic explainers such as LIME (Ribeiro et al. 2016) and SHAP (Lundberg and Lee 2017). These methods check how a metric’s sentence level score changes when the input sentence is perturbed at selected positions. While the results correlate with the human annotations to some degree, techniques that use model access performed better in the shared tasks. Gradient based methods are explored by Fomicheva et al. (2022); Eksi et al. (2021); Treviso et al. (2021); Rei et al. (2022a). These papers use methods like Integrated Gradients (Sundararajan et al. 2017) to determine the effect each input token has on the output. Attention values perform especially good for Treviso et al. (2021); Rei et al. (2022b), leading to winning techniques for most language pairs in both shared tasks, i.e., producing the most plausible explanations. In specific, Treviso et al. (2021) succeed by scaling attention weights by the norm of value vectors and filtering for attention heads that have the highest scores in detecting errors. Rei et al. (2022b) build on their approach and (1) instead scale attention weights with the l2-norm of the gradient of value vectors and (2) designing an ensemble method for attention heads. In their final submission, they ensemble the best performing attention heads over all layers. Word-level contextualized embeddings are used by Leiter (2021); Tao et al. (2022); Azadi et al. (2022). Specifically, they leverage the word-level embedding distances that are inherently computed in token-matching metrics (see §2.1) and use them as importance scores. For model-specific works, Rubino et al. (2021) jointly fine-tune their metrics on the word-level and sentence level using synthetic data, to directly output the computed word-level scores.131313Rubino et al. (2021) won the constrained track of Eval4NLP21, which did not allow for word-level finetuning. Kabir and Carpuat (2021) use LIME with Transquest (Ranasinghe et al. 2020a) and merge the word-level importance scores with additional word-level scores obtained from a pure word-level model. After the shared tasks, Fernandes et al. (2022b) enhance explainability techniques (in case of MT they use a gradient technique) by adding learnable parameters and employing a student teacher mechanism to improve the correlation to human word-level scores. Rei et al. (2023) extend the work of Rei et al. (2022b) by using reference-based metrics (e.g. COMET) and importance scores of the references. They construct a data set from the WMT21 MQM annotations Freitag et al. (2021b) by assuming that feature attribution scores link to annotated error spans. Evaluations on this data set show that the winning explainability techniques of the shared task perform best and that references can improve the plausibility. Notably, they also show that explanations can help to detect critical errors in translations. Based on this, they employ the explanations for metric diagnosis, to test whether their metrics are more prone to ignoring certain types of critical errors. In their limitations, they note that the explainabiltiy techniques they use have shown to be faithful in some scenarios, but also note that they did not specifically test whether they are faithful for their architecture. Finally, we note that many more metrics can provide feature-importance like explanations by design. Similar to the base metrics for Leiter (2021); Tao et al. (2022); Azadi et al. (2022), most unsupervised matching metrics could be used to determine importance scores based on word- or embedding-similarities. Further, the per-word generation probabilities in text generation metrics like PRISM or BARTScore can also be interpreted as feature importance (they can also be interpreted as the confidence of the generation models; Fomicheva et al. 2020).
In the example presented in Figure 5, high feature importance scores are assigned to words that were translated correctly, i.e. “Things/complicated” and lower scores indicate words that are wrong.
Leiter et al. (2022b) take a different approach on using feature importance techniques for the goal of metric improvement. Instead of using the feature importance scores in a task that might involve human explainees, the authors aggregate and directly integrate the feature importance scores into metric scores. This follows the intuition that the explanations carry additional information that can be beneficial to the sentence-level correlation of a metric with human judgments.
Goals: The discussed feature importance approaches are tested for plausibility, so strong approaches could be used for metric accessibility/improvement and semi-automatic labeling. The usage for metric diagnosis is currently limited, as only samples are explained, and approaches have not been tested for faithfulness. The current approaches have not been used to detect social biases.
Audiences: The relation of low/high scores and translation errors is intuitive, such that these techniques could be used by non-expert and expert users. Metric developers could use the methods to create new data sets or to directly incorporate them into metrics Leiter et al. (2022b).
Limitations: For all approaches, it is not specifically tested whether the explanations are faithful.141414In different possible aspects (Jacovi and Goldberg 2020) Notably, Rei et al. (2023) consider that faithfulness is probably given in their case but note the missing tests as limitation.
Second, certain translation errors cannot be easily captured by highlighting specific words. For example, the word-level scores cannot represent cases where the MT fails to explicitly express a grammatical feature that is omitted in the source language but is required to be explicit in the target language (for example, the use of articles when translating from Russian into German).
Third, the explanation type does not provide correspondence between highlighted error words in the source and target language.
Fine-grained errors (Row 3 in Figure 4) Some recent works detect errors on a word-level and leverage their number and type to compute a sentence-level score (Lu et al. 2022, 2023; Bao et al. 2023; Xu et al. 2023). This follows the ideas of the MQM annotation scheme (Lommel et al. 2014; Freitag et al. 2021a) in which human experts evaluate translations in a similar manner. Freitag et al. (2021a) have shown that this annotation scheme improves over previous annotation methods such as direct assessments (Graham et al. 2015). While the main goal of these works is to provide sentence-level scores, they are also explainable-by-design and the generated error-labels carry explanatory value (model-specific in Figure 4). In specific, BARTScore++ (Lu et al. 2022) distinguishes explicit errors that are easy to detect and implicit errors. They detect the explicit errors based on BART generation probabilities. Then they generate a sentence-level score based on an additional improved hypothesis sentence in multiple BARTScore calculations. Lu et al. (2023) also combine fine-grained errors into a score. They first request ChatGPT to generate explanations based on MQM annotations (erroneous words, error type and error majority). In a second step, they ask a model to calculate a score from these. They note the instabilities of ChatGPT, but find that their approach is strong on a system-level and generates sentence-level scores that are better than BERTScore. Xu et al. (2023) fine-tune InstructScore, a LLaMA model (Touvron et al. 2023) that generates a text containing error majority, error spans, error types (MQM) and an explanation for every error in an input sentence. The generated texts are similar to Lu et al. (2023) but add more details on each error. Similarly, they construct a score from the count and severity of errors. They obtain synthetic training data from GPT4 (OpenAI 2023a), analyze failure modes of InstructScore during training and later optimize the metric on real machine translation outputs on data sets with annotated ground truth scores. Their model shows state-of-the-art results for reference-free sentence-level evaluations. Further, the detailed error annotations can be used as fine-grained explanations. Differently to the previous approaches, Bao et al. (2023) fine-tune a seq2seq model to predict addition and omission errors on a word-level. In the example presented in Figure 5, a tag “sense” flags the wrong word sense of “became”. Notably, all fine-grained error approaches we describe are very recent, which is potentially caused by the adaptation of MQM in the WMT metrics shared tasks and the recent successes of LLMs.
Goals: These fine-grained error approaches can be used for increased accessibility and semi-supervised labeling, as the methods model plausible explanations (some of them by testing the alignment of scores aggregated from explanations with human MQM scores). The usage for metric diagnosis is limited, as only samples are explained, and explanations have not been tested for faithfulness. The current approaches cannot be used to detect social biases.
Audiences: Flagged translation errors are simple to understand, such that these techniques could be used by non-expert and expert users. Metric developers might use them to create new data sets to train new metrics.
Limitations: The current approaches (besides Xu et al. 2023) do not have the goal to provide explanations that make the sentence-level scores more descriptive. Instead, they only search to build sentence-level metrics from LLM
outputs. As such, there might be open opportunities in leveraging the explanations provided by these models to explain to a human explainee. Another limitation are missing evaluations of faithfulness.
Perturbation robustness (Row 4 in Figure 4) The methods described up to now addressed decision-understanding techniques Gehrmann et al. (2020). This paragraph and the next describe model-understanding techniques (we describe these terms in §2.2). Many recent works evaluate the robustness of MT metrics (Sai et al. 2021; Karpinska et al. 2022; He et al. 2022; Sun et al. 2022; Chen and Eger 2023; Vu et al. 2022). Robustness evaluation qualifies as model-understanding, as it usually infers properties that describe a model (see §2.2). All works in this category are structured into “black box” in Figure 4.
Sai et al. (2021) propose perturbation checklists, following Ribeiro et al. (2020), that evaluate how susceptible metrics are towards predefined types of perturbations as evaluation criterion. This allows developers to check whether all invariances that are required for a specific task are fulfilled. In particular, they compare the change in the metric’s score with the change in score a human would assign after the perturbation. These templates encompass dropping or adding of context as well as negations. They show that perturbation checklists allow to pick metrics that are strong with respect to specific properties. For example, their experiments indicate MoverScore (Zhao et al. 2019) would capture fluency better than BERTScore (Zhang et al. 2020a) due to its ability to cope with the jumbling of words. The checklists allow for more fine-grained assessment of automatic evaluation metrics, which exposes their limitations and supports the design of better metrics. Chen and Eger (2023) create a similar, but fully automatic variant of this evaluation. They design several perturbation types (framed as adversarial attacks) such as addition, omission and negation. Then they measure whether a tested metric prefers a meaning preserving, but lexically dissimilar paraphrase of a source/reference sentence over a minimally attacked but lexically similar version of the source/reference sentence. This means they check whether the metric score of the attacked example is successfully lower than the metric score of a paraphrase. One of their findings is that standard metrics often have problems in detecting wrong names or numbers in translations. Karpinska et al. (2022) create DEMETR, a data set of machine translations that were perturbed using several perturbation types. They propose to test a metric’s preference between these perturbations and the original translation. Their task is easier for a metric to compute than Chen and Eger (2023), as the preference is checked directly between attacked and non-attacked version, while Chen and Eger (2023) use a dissimilar paraphrase instead of the non-attacked version. Also, they define more perturbation types than Chen and Eger (2023), but require manual annotation for some of them. He et al. (2022) also compare the original translation with certain perturbation types. They differ from the other two papers in terms of domains they are applying their method to and in the perturbation types that they propose. Vu et al. (2022) use character-level adversarial attacks to evaluate the robustness of several BERTScore variants.
The WMT21 (Freitag et al. 2021b) and WMT22 (Freitag et al. 2022) metrics shared tasks introduce a similar setting for robustness tests. In WMT21, the organizers provided the participants with a challenge set composed of translations that were perturbed in a specific way, to test whether metrics reduce their score as expected when presented with respective phenomena. In WMT22, participants of a challenge set subtask (Alves et al. 2022; Avramidis and Macketanz 2022; Chen et al. 2022; Amrhein et al. 2022) designed data sets posing challenges in a similar manner. In the example presented in Figure 5, we show tests similar to those performed by techniques of this category. The metric score should be decreased if numbers or word order used in the hypothesis change.
The work by Guerreiro et al. (2023) is also related to these approaches. One of their contributions is an expert annotated data set of machine translation hallucinations. They use this data set to test how well selected metrics can identify each hallucination type. Further, they give the recommendation to simply use the sequence log-probability for hallucination detection.
While the previous techniques describe the susceptibility of metrics towards various perturbation types, Sun et al. (2022) describe methods to analyze the metrics’ social biases. They also base their evaluation on perturbed input sentences; however, they evaluate whether a metric grades two versions of a sentence that contain opposite social stereotypes the same with respect to a neutral reference sentence (that does not contain stereotypes). They find that generation-based metrics show the least biases.
Goals: The presented perturbation robustness approaches give detailed insights into metric weaknesses. Therefore, they are most suited to be used for metric diagnosis and improvement.
Audiences: Metric developers can use these methods for diagnosis. They can also be used by metric users to check if a metric fits their use case. Regulatory entities might test if metrics fulfill their guidelines, e.g. exhibit no social biases. Lastly, affected users might gather evidence when treated unfairly.
Limitations: A drawback of the perturbation-based approaches is that they require predefined error types. Therefore, errors lacking a definition in the evaluation scheme that would fool the machine translation metrics could be missed. These approaches are faithful to some degree, namely in the domains they have been evaluated in. For small perturbation sets it could be questionable if the model generally behaves as predicted. Therefore we would suggest evaluating each property on a large, general data sample. Still, it could be advisable to employ faithfulness tests, if possible, to achieve guarantees.
Linguistic properties (Row 5 in Figure 4) Here, we summarize model-understanding techniques that directly summarize various linguistic properties a metric fulfills. So far, the only method that falls into this category for MT is Kaster et al. (2021) (see linguistic properties/black box in Figure 4). They propose a model-understanding technique that decomposes the score of sentence-level BERT-based metrics into linguistic factors. In particular, they explore the degree to which metrics consider the properties syntax, semantics, morphology and lexical overlap. To do so, they measure these properties in a given data set and learn regressors that explain how much each distinct property contributes to the original metric’s score. In their experiments, they show that each metric captures semantic similarity and lexical overlap to some degree. Syntactic and morphological similarity are either captured to a smaller extent or are even negatively correlated with the metric score. Especially, MoverScore (Zhao et al. 2019) and BERTScore (Zhang et al. 2020a) have a comparatively high coefficient for the lexical score.
As shown in Figure 5, these techniques might give explanations like “the metric puts a high focus on lexical properties (green bar) and a low focus on semantics (red bar)“.
Note that the explanation of linguistic properties is close to the decision-understanding of quality aspects. If these are collected over many samples and a regression is performed, similar explanations might be achieved. A related approach used for another NLP task is by Opitz and Frank (2022), who retrain sentence embeddings to decompose into sub-components that match the aspects of abstract meaning representation, a format to represent the meaning of sentences. This allows them to attribute parts of the score to different aspects of a sentence. They evaluate their approach on the model-understanding level. When this would be applied to MT, it would also count towards this explanation type.
We also list Kaster et al. (2021); Opitz and Frank (2022) with quality aspect explanations, as the metrics that measure the linguistic factors can also be computed on a sentence level.
Goals: The discussed linguistic properties approaches evaluate general metric properties and can be used for metric debugging. As they generate sentence/metric-level property scores, the techniques
are
not usable for word-level semi-automatic labeling. When used on a sentence-level, the techniques can aid the goal of making metrics more accessible. While not employed here, the techniques might be usable to detect social biases by introducing a set of bias detection scores from which to learn the regressors.
Audiences: The linguistic properties can be easily explained to non-experts, however metric selection should not solely be performed based on linguistic properties
and
is advised to be handled by domain experts. The methods are especially useful to debug metrics, i.e. for metric developers.
Limitations: The approach by Kaster et al. (2021) could not explain reference-free metrics well, so plausibly requires alternate explanatory variables.
The search for regressors may be inspired by quality aspect approaches (described earlier this section) where not a global metric score is reported but several scores (such as coherence, fluency, etc.). These could then compose a global MT score.
Further, some of the “property metrics” could be considered black box variables themselves and future work might replace them by more transparent factors.
One might also explore the collinearity of the different regressors.
For Kaster et al. (2021), faithfulness is likely given to some degree, around the samples that was trained on. Still, it could be advisable to employ faithfulness tests, if possible, to achieve guarantees.
5 Future Work
[set/.style = draw]
[…]
“Die Dinge wurden kompliziert”Score:1
“Die Dinge erhalten kompliziert”Score:0.5;
[…]
In German, the verb ”bekommen” means ”to get” or
”to receive,” but in this context, it doesn’t convey the
intended meaning effectively. Instead, we can use the
verb ”werden” which means ”to become” to better
capture the sense of change or transformation.;
[…]
“became” Source & “werden” Hypothesis
Score;
[…]
Importance((“became”,“bekommen”)): 0.1
Importance((“Things”,“Die Dinge”)): 0.6;
[…]
Neuron 1423: bekommen
Neuron 5234: werden;
[…]
Explanation Types;
[…]
Ground Truth: “Things became complicated”
Hypothesis: “Die Dinge bekamen kompliziert”
Score: ;
[[sep=6.2pt]-Computer Modern Rightarrow[length=3mm, width=3mm], line width=0.7mm, shorten ¡=8pt, shorten ¿=8pt] (A) edge (F) (B) edge (F) (C) edge (F) (D) edge (F) (E) edge (F) ;
In this section, we describe future research directions for explainable MT metrics. Specifically, we (1) consider directions for explanation types that have been used in the papers described in our taxonomy, (2) describe potential explanation types from other domains, (3) point out needs for future evaluations and concepts and (4) urge for an explanation of domain boundaries.
Exploring present explanation types Based on our taxonomy (see Figure 4), we identify unexplored areas of explainable MT evaluation (empty table cells). We find that only feature importance techniques are thoroughly explored on all levels of model access. The other methods mostly explain in black box settings or are based on metrics that are explainable by design. We suppose that gradients, attention weights and embeddings have so far mostly been applied in the field of feature importance, as there is often already a 1-1 relationship their values have towards input features. Also, due to the two recent shared tasks, feature importance has likely received the most attention.
Future work could consider enhancing quality aspect explanations — Row 1 in Figure 4 — by incorporating more model access. For example, a separate fluency score might be realizable by employing measurements on target side embeddings in unsupervised matching metrics. Current model-understanding techniques for MT metrics — Row 4-5 in Figure 4 — are mostly applied in a black box manner. For these, using more model access might also provide further insights. For example, analyses of the contextualized embedding space and how it affects a metric score could explain the metric on a more general level. Many works consider the probing of transformer-based architectures (e.g. Tenney et al. 2019; Liu et al. 2019; Sajjad et al. 2022b) and it would be interesting to see if their results can inform the metric development.
Also, current perturbation robustness techniques often use manually defined perturbation strategies. It would be interesting to use adversarial attacks instead that automatically infer these strategies.
Exploring other explanation types Besides the works in our taxonomy, other fields of NLP and explainable AI in general have proposed further explanation types, e.g. counterfactuals, natural language explanations and rule extraction (e.g. Madsen et al. 2022). We illustrate these options with an example in Figure 6. We already introduced counterfactuals in §2.2, stating their role in evaluations of robustness (as adversarial attacks) and fairness. It remains to explore if single sentence explanations also carry explanatory value. In the examples for counterfactuals in Figure 6, the original word “bekommen” is switched with the words “wurden” (to get) resp. “erhalten” (to receive) in two counterfactuals. Based on the scores achieved by these counterfactuals, a user could infer metric weaknesses. One technique to generate counterfactuals is for example described by Wu et al. (2021).
Decision rule explanations describe rules which a model follows when computing its output Ribeiro et al. (2018). Often, these techniques were designed to explain classification models and are based on perturbations that change the originally predicted class Ribeiro et al. (2018). For metrics, the scores could be discretized beforehand. An example from other NLP domains are Anchors Ribeiro et al. (2018). In our example (see Figure 6), the rule describes that if “became” is present in the source and “wurden” is present in the translation, the metric will assign a score higher than 0.5. Obviously in other contexts, such a rule would fail.
Another type of explanations are hierarchical feature importance explanations, which will assign importance scores to groups of input tokens in single Chen et al. (2020) or paired input sentences Chen et al. (2021). In the given hypothetical example (see Figure 6), the groups of became/bekommen and Things/Die Dinge were formed and get assigned separate importance scores. A related approach was tested in the SemEval-2015 and -2016 shared tasks on semantic similarity, where the task was to assign labels that capture the similarity between token groups from two sentences Agirre et al. (2015, 2016).
Natural language explanations are explanations in free text form (e.g. Wiegreffe and Marasovic 2021). This is an explainability type where the generative abilities of recent large language models (LLM), like ChatGPT and GPT4, can be leveraged. The works by Fu et al. (2023); Lu et al. (2023); Kocmi and Federmann (2023) already explore the use of LLMs as metrics and extract the metric scores from LLM outputs (or generation probabilities) that contain (or could easily be extended to provide) natural language descriptions of why a certain score is assigned. In contrast, Xu et al. (2023) explicitly highlight the explainability aspect of their model. In our taxonomy, we grouped them under fine-grained error explanations, as this is their main goal. However, they also produce natural language explanations for each error.151515We again note that these techniques will not necessarily describe a metric’s internal workings, but might be usable due to providing plausible explanations, see §2.2. In these works, the explanations are rather used as a byproduct, while an evaluation of the explanation quality and further use cases remains for future work.
The example in Figure 6 was generated with the help of ChatGPT, which shows that such an approach might yield plausible explanations. An example from other NLP domains is Rajani et al. (2019), who generate explanations for commonsense reasoning. In NLG in general, many other approaches for LLM-based metrics have been constructed since the release of ChatGPT. These could potentially be extended to MT metrics and explainability (e.g. Chiang and yi Lee 2023; Wang et al. 2023; Ji et al. 2023a; Chen et al. 2023; Liu et al. 2023).
Recent works, describe prompting strategies like chain-of-thought (Wei et al. 2022) and tree-of-thought (Yao et al. 2023) where language models show performance improvements if they are giving explanations of their output. It could be an interesting research direction, to see if the explainability techniques summarized in this paper can improve metric performance in this way, or if these techniques can provide reasonable explanations (that are at least plausible).
Neuron Explanations are another potential use case of LLMs to separately explain each Neuron in the underlying networks. Bills et al. (2023) recently showed that GPT4 can (to some extent) summarize highlighted neuron activation in textual input of GPT2 to the same extent humans can. This allows to successfully simulate the behavior and effects of some neurons. For MT metrics, this would allow to determine the role certain neurons play to determine scores when specific features like “became is used” are present in the metric input. At the moment, these methods however still have some limitations Bills et al. (2023).
Interactive Explanations Due to their context size, instruction fine-tuning and emergent abilities of recent LLMs allow to interactively ask questions and adapt based on the answers (e.g. OpenAI 2023b). Jacovi et al. (2023) state that interactivity is a key point of successfully communicating the explained subject to an explainee. LLM-based explanations (like natural language/neuron explanations …) might achieve this. For example, a user could ask for more detailed information for a mistranslation that was noted in a first explanation.
New approaches could furth consider displaying multiple explanation types together, like Xu et al. (2023) who present fine-grained error annotations with natural language explanations.
Future evaluation and concepts Current work scarcely performs human evaluation of explanations. This is partly because the focus is often not on the explanations but other aspects, like the sentence-level scores. One form of human evaluation (for plausibility) is often to use pre-computed human explanations and check the correlation to them (e.g. Zerva et al. 2022). Current works on MT metrics do, however, not evaluate whether the explanations are actually helpful to explain the concepts they describe to their audience. This means, whether the explanations manage to help the explainees to adapt their own theory of mind to be a coherent abstraction of the real problem Jacovi et al. (2023). Hence, we encourage future work to address this, e.g. by considering simulatability Hase and Bansal (2020). Regarding plausibility, it could also be interesting to evaluate whether plausible explanations of incorrect metric scores can actually deceive non-experts or experts (with the current state-of-the-art), as warned by Jacovi and Goldberg (2021); Jin et al. (2023).
As noted throughout the paper, the degree of faithfulness an explanation has to the metric’s internal decision process is an important property (Jacovi and Goldberg 2020), especially for metric debugging. Current works on decision-understanding of MT metrics, however, mostly dismiss faithfulness. Future work might consider the evaluation of metric faithfulness, to enable these use-cases. We note that this evaluation will be complex, due to faithfulness being multi-faceted (Jacovi and Goldberg 2020). For example, lower-granularity scores can faithfully explain sentence-level scores that are aggregated from them, however this does not explain how the actual black box component produced the lower-granularity scores. Further, there are pitfalls that exist in faithfulness evaluation (Ju et al. 2022), warning against the usage of common methods like area over the perturbation curve scores (AOPC) (Samek et al. 2017). Therefore, future work should first consider what constitutes faithful explanations in terms of metrics and design evaluation approaches.
Further, we note that current works mostly consider explaining sentence-level scores. Future work, might also consider other metric granularity types more.
Exploring domain boundaries Future work might also consider under-explored use cases of explainability for MT metrics. Currently, only one work considers the social biases of MT metrics Sun et al. (2022) and none considers the usage of explainability for semi-automatic labeling of training data. Besides examining just the metrics, future work could also inspect the relation of explainable MT metrics, especially plausible-only explanations, to explainable MT (e.g. Stahlberg et al. 2018), as explanations of both will probably be similar. Lastly, future work could explore providing similar taxonomies as ours that are comprising the whole field of explainable natural language generation metrics.
6 Related Work
A large number of conceptual works and surveys have been conducted in the area of explainability (e.g. Lipton 2016; Biran and Cotton 2017; Adadi and Berrada 2018; Doshi-Velez and Kim 2017; Došilović et al. 2018; Guidotti et al. 2018; Arya et al. 2019; Carvalho et al. 2019; Gilpin et al. 2019; Miller 2019; Barredo Arrieta et al. 2020; Bodria et al. 2023; Linardatos et al. 2021).
There are few surveys that touch upon the need for explainable NLG (here MT), thereby motivating our work. In their survey of NLG metrics, Celikyilmaz et al. (2020) see the need for explainable evaluation metrics to spot system quality issues and to achieve higher trust in the evaluation of NLG systems. They consider that such quality issues might be unintended biases and factual inconsistencies. Sai et al. (2022) instead propose that explainable NLG metrics should focus on providing more information than just a single score (such as fluency or adequacy). Gehrmann et al. (2023) also notice the need for interpreting NLG metrics and give a brief overview of selected related approaches. The main contribution of their work is an overview of common NLG metric shortcomings and best practices to prevent them. Our focus on concepts for explainable MT metrics is a different one and our taxonomy of approaches is much more extensive.
General surveys on explainable NLP are related to our work, as they highlight how explainability techniques have been used in NLP and might further guide the development of future approaches for explainable MT metrics (e.g. Vijayakumar 2023; Sajjad et al. 2022a; Gurrapu et al. 2023; Lertvittayakumjorn and Toni 2021; Saha et al. 2022; Sun et al. 2021; Zini and Awad 2022; Qian et al. 2021; Madsen et al. 2022). Besides these works, there are also surveys of robustness (e.g. Goyal et al. 2023) and fairness (e.g. Blodgett et al. 2020), which are related, as discussed in §2.2.
The taxonomy graphic in Figure 4 was adapted from the survey of post-hoc explainability methods in NLP by Madsen et al. (2022). In specific, we follow their approach to display model access on the x-axis and explanation types on the y-axis. Further, we have customized their TIKZ code. We specifically focus on the explainability of MT metrics; therefore, our taxonomy is completely different content-wise. Our definitions of explainabilty, its goals and audiences are in some parts adapted from the survey of Barredo Arrieta et al. (2020) (see §2.2). Our work is different from previous surveys, as we set our focus on the explainability of MT metrics.
7 Conclusion
In this work, we discuss audiences, goals and properties for explainable machine translation metrics, a nascent field that may help further overcome the dominance of classical low-quality evaluation metrics. We also survey and categorize recent approaches on explainable MT metrics into a taxonomy, highlighting their results and limitations. Currently, two dominant approaches to explainability for MT metrics are (1) feature importance explanations that highlight erroneous words in source and hypothesis to explain sentence-level scores and (2) perturbation robustness approaches that check a metric’s robustness to manually devised types of input perturbations. We also identify a current trend towards (3) fine-grained error explanations likely caused by the recent improvements of LLMs like ChatGPT and by the adaptation of MQM annotations in the WMT metrics shared tasks. A major weakness of the current realization of (1) is that error highlights do not consider the correspondence between words in source and target. Further, a weakness of (2) is that perturbation types have to be manually defined, an issue that might be tackled by using more general approaches of adversarial attacks in the future (we discuss this issue in more detail in an earlier version of this paper; Leiter et al. 2022a). Generally, most current evaluations of explainability approaches for MT metrics do not consider the faithfulness of explanations, limiting their use in metric debugging and high-risk scenarios. We also present a vision of next-generation approaches to explainable evaluation metrics, which should help fix the problems of the above paradigms and provide guidance to explore unexplored areas. Here, we also urge future work to consider the desiderata and implications of faithfulness for explainable MT metrics.
Our broader vision is that explainability is now a ‘desirable but optional’ feature, but we argue that in the future it will become essential, even compulsory, especially for evaluation metrics as a highly sensitive task assessing the quality (and veracity) of translated information content. Explainability builds transparency and trust for users, eases bug-fixing and shortens improvement cycles for metric designers and will be required by law/regulations for AI systems to be applied to large-scale, high-stake domains. In this context, we hope our work will catalyze efforts on the topic of explainable evaluation metrics for machine translation.
Acknowledgements
Since November 2022 Christoph Leiter is financed by the BMBF project “Metrics4NLG”. Piyawat Lertvittayakumjorn had been financially supported by Anandamahidol Foundation, Thailand, from 2015-2021. He mainly contributed to this work only until September 2022 while affiliated with Imperial College London (before joining Google as a research scientist after that). Marina Fomicheva mainly contributed to this work until April 2022. Wei Zhao is supported by the Klaus Tschira Foundation, Heidelberg, Germany. Yang Gao mainly contributed to this work before he joined Google Research in December 2021. Steffen Eger is financed by DFG Heisenberg grant EG 375/5–1 and by the BMBF propject “Metrics4NLG”.

### v2

@@
-shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation
+Towards Explainable Evaluation Metrics for Machine Translation
 Abstract
@@
 We hope that our work can help catalyze and guide future research on explainable evaluation metrics and, mediately, also contribute to better and more transparent machine translation systems.
+Keywords: evaluation metrics, explainability, interpretability, machine translation, machine translation evaluation
 1 Introduction
-The field of evaluation metrics for Natural Language Generation (NLG), especially machine translation (MT) is in a crisis (Marie et al.
-2021).111We published an earlier version of this paper (Leiter et al.
-2022a) under a different title.
-Both versions consider the conceptualization of explainable metrics and are overall similar.
-However, while the old version contains several novel experiments, this new version puts a stronger emphasis on the survey of approaches for the explanation of MT metrics including the latest LLM based approaches.
-For example, this comprises techniques that return fine-grained error labels or natural language explanations.
-Despite the development of multiple high-quality evaluation metrics in recent years (Zhao et al.
-2019; Zhang et al.
-2020a; Rei et al.
-2020; Sellam et al.
-2020; Yuan et al.
-2021), the Natural Language Processing (NLP) community appears hesitant to adopt them for assessing NLG systems (Marie et al.
-2021; Gehrmann et al.
-2023).
+The field of evaluation metrics for Natural Language Generation (NLG), especially machine translation (MT) is in a crisis (Marie et al., 2021).
+Despite the development of multiple high-quality evaluation metrics in recent years (e.g., Zhao et al., 2019; Zhang et al., 2020a; Rei et al., 2020; Sellam et al., 2020; Yuan et al., 2021; Rei et al., 2023a; Kocmi and Federmann, 2023a), the Natural Language Processing (NLP) community appears hesitant to adopt them for assessing NLG systems (Marie et al., 2021; Gehrmann et al., 2023).
 Empirical investigations of Marie et al.
-(2021) indicate that the majority of MT papers relies on surface-level evaluation metrics such as BLEU and METEOR (Papineni et al.
-2002; Banerjee and Lavie 2005), which were created two decades ago, a trend that may allegedly have worsened in recent times.
-These surface-level metrics cannot (even) measure semantic similarity of their inputs and are thus fundamentally flawed, particularly when it comes to assessing the quality of recent state-of-the-art MT systems (e.g.
-Peyrard 2019; Freitag et al.
-2022), raising concerns about the credibility of the scientific field.
+(2021) indicate that the majority of MT papers relies on surface-level evaluation metrics such as BLEU and METEOR (Papineni et al., 2002; Banerjee and Lavie, 2005), which were created two decades ago, a trend that may allegedly have worsened in recent times.
+These surface-level metrics cannot (even) measure semantic similarity of their inputs and are thus fundamentally flawed, particularly when it comes to assessing the quality of recent state-of-the-art MT systems (e.g., Peyrard, 2019; Freitag et al., 2022), raising concerns about the credibility of the scientific field.
 We argue that the potential reasons for this neglect of recent high-quality metrics include: (i) non-enforcement by reviewers; (ii) easier comparison to previous research, for example, by copying BLEU-based results from tables of related work (potentially a pitfall in itself); (iii) computational inefficiency to run expensive new metrics at large scale; (iv) lack of trust in and transparency of high-quality black box metrics.
@@
 In recent years, explainability has become a crucial research area in AI due to its potential benefits for users, designers, and developers of AI systems Samek et al.
-(2018); Vaughan and Wallach (2021).222 As of now, there is no universally accepted definition of explainability in the AI community.
+(2018); Vaughan and Wallach (2021).111As of now, there is no universally accepted definition of explainability in the AI community.
 In this work, we adapt the definition by Barredo Arrieta et al.
@@
 (2020) stress the need for explainable evaluation metrics to spot system quality issues and to achieve a higher trust in the evaluation of NLG systems.
-Explanations indeed play a vital role in building trust for new evaluation metrics.333As an illustrating example, Moosavi et al.
+Explanations indeed play a vital role in building trust for new evaluation metrics.222As an illustrating example, Moosavi et al.
 (2021) distrust using the metric BERTScore Zhang et al.
@@
 In fact, BERTScore typically has a particularly narrow range, so a score of 0.74 even for bad outputs may not be surprising (BERTScore provides a scaling method that solves these issues to some degree).
-Explainability techniques would be very helpful in preventing such misunderstandings, e.g.
-by showing that the features BERTScore attends to align with human judgement.
-For instance, one potential reason that a metric will likely be better accepted by the research community is if the explanations of the scores align well with human reasoning and faithfully reflect its internal decision process (see §2.2).
+Explainability techniques would be very helpful in preventing such misunderstandings, e.g., by showing that the features BERTScore attends to align with human judgement.
+For instance, if explanations of the scores align closely with human reasoning and faithfully reflect a metric’s internal decision process, this metric may be more likely to gain acceptance within the research community (see §2.2 for a definition of faithfulness and related terminology).
 By contrast, if faithfulness is given and the explanations are counter-intuitive, users and developers might lower their trust and be alerted to take additional actions, such as trying to improve the metrics using insights from the explanations or looking for alternative metrics that are more trustworthy.
-Furthermore, explainable metrics can be used for other purposes: for example, when a metric produces a low score for a given input, highlighted words (a widely used method for explanation, see §4) in the input are natural candidates for manual post-editing.
+Furthermore, explainable metrics can be used for other purposes: for example, when a metric produces a low score for a given input, highlighted words (a widely used method for explanation, see §4) in the input are natural candidates for manual post-editing (when working with reference-free metrics, see §2.1).
 This concept paper aims at providing a systematic overview of the existing efforts in explainable MT evaluation metrics and an outlook for promising future research directions.
-The main focus of this work lies on explainable evaluation metrics for MT, many of our observations and discussions can likely be adapted to other NLG tasks, however.
+The main focus of this work lies on explainable evaluation metrics for MT, although many of our observations and discussions can likely be adapted to other NLG tasks.
 We make the following contributions, outlined by the structure of our paper:
@@
 §5 Future Work: An analysis of underexplored research directions, and a collection of exemplary methods from other NLP domains to illustrate potential future paths.
-Here, we also discuss the usage of large language models (LLMs), like ChatGPT OpenAI (2023b) and GPT4 OpenAI (2023a).
-With this work, we aim to solidify the field of explainable MT metrics and provide guidance for researchers, in specific metric developers, who aim to better understand and explain MT metrics, as well as users that want to employ explainability techniques to explain metric outputs.
+Here, we also discuss the usage of recent large language models (LLMs), like ChatGPT OpenAI (2023a) and GPT4 OpenAI (2023b).
+With this work, we aim to solidify the field of explainable MT metrics and provide guidance for researchers, e.g., metric developers, who aim to better understand and explain MT metrics, as well as users that want to employ explainability techniques to explain metric outputs.
 In the long term, we envision that our work will aid the development of improved MT metrics and thereby help to improve the quality of machine translations.
@@
 MT metrics grade machine translated content, the hypothesis, based on ground truth data.
-We differentiate these metrics along several dimensions.
+Metrics can be categorized along several dimensions (e.g., Sai et al., 2022; Celikyilmaz et al., 2020).
+Here, we focus on a minimal specification that we will leverage later.
 A summary is shown in Table 1.
-We note that metrics can be categorized along many dimensions (e.g.
-Sai et al.
-2022; Celikyilmaz et al.
-2020).
-Here, we focus on a minimal specification that we will leverage later.
 Some parts of this section refer to explainability (see definitions in §2.2).
 Dimension | Description |
-| Input type |
-Whether source, reference translation or both are used as ground truth for comparison |
-| Granularity |
-At which level a metric operates: Word-level, sentence-level, document-level |
-| Quality aspect |
-What a metric measures: Adequacy, fluency, etc. |
-| Learning objective |
-How a metric is induced: Regression, ranking, etc. |
-Input type We call a metric reference-based if it requires one or multiple human reference translations to compare with the hypothesis as ground truth, which can be seen as a form of supervision (e.g.
-Papineni et al.
-2002; Zhao et al.
-2019; Rei et al.
-2022a).
-Reference-free metrics do not require a reference translation to grade the hypothesis (e.g.
-Zhao et al.
-2020; Ranasinghe et al.
-2020b; Belouadi and Eger 2023).
+Input type | Whether source, reference translation or both are used as ground truth for comparison |
+Granularity | At which level a metric operates: Word-level, sentence-level, document-level |
+Quality aspect | What a metric measures: Adequacy, fluency, etc. |
+Learning objective | How a metric is induced: Regression, ranking, etc. |
+Input type We call a metric reference-based if it requires one or multiple human reference translations to compare with the hypothesis as ground truth, which can be seen as a form of supervision (e.g., Papineni et al., 2002; Zhao et al., 2019; Rei et al., 2022a).
+Reference-free metrics do not require a reference translation to grade the hypothesis (e.g., Zhao et al., 2020; Ranasinghe et al., 2020a; Belouadi and Eger, 2023).
 Instead, they directly compare the source to the hypothesis.
-In the literature, reference-free MT evaluation is sometimes also referred to as “reference-less” (e.g.
-Mathur et al.
-2020) or “quality estimation” (e.g.
-Zerva et al.
-2022).
+In the literature, reference-free MT evaluation is sometimes also referred to as “reference-less” (e.g., Mathur et al., 2020) or “quality estimation” (e.g., Zerva et al., 2022).
 This dimension is important for explainable MT evaluation, as explainability techniques might need to consider which resources are available.
 Granularity Translation quality can be evaluated at different levels of granularity: word-level, sentence-level and document-level.
-The majority of metrics for MT return a single sentence-level score for each input sentence (e.g.
-Zhao et al.
-2019; Zhang et al.
-2020a).
-Beyond individual sentences, a metric may also score whole documents (multiple sentences) (Jiang et al.
-2022; Zhao et al.
-2023).
-In the MT community, metrics that evaluate translations at the word-level (for example, whether individual words are correct or not) are also common (Turchi et al.
-2014; Shenoy et al.
-2021).
-Recently, there has been a tendency to compare metrics to human scores derived from fine-grained error annotations — multidimensional quality metrics (MQM) (Lommel et al.
-2014) — as they tend to correspond better to human professional translators (Freitag et al.
-2021a).
-Metrics of higher granularity (e.g.
-word-level) provide more explanatory value than those of lower granularity, as they provide more details on which parts of the input might be translated incorrectly (e.g.
-Leiter 2021; Fomicheva et al.
-2021, 2022).
-Most techniques we describe in §4 target the explainability of sentence-level metrics.
-Quality aspect This refers to the properties a metric evaluates and measures, for example, human assigned adequacy (via direct assessment (Graham et al.
-2016) obtained from crowd-workers), fluency, or other aspects (such as relevance, informativeness, or correctness, mostly in other NLG fields) (e.g.
-Sai et al.
-2022; Celikyilmaz et al.
-2020).
+The majority of metrics for MT return a single sentence-level score for each input sentence (e.g., Zhao et al., 2019; Zhang et al., 2020a).
+Beyond individual sentences, a metric may also score whole documents (multiple sentences) (Jiang et al., 2022; Zhao et al., 2023).
+In the MT community, metrics that evaluate translations at the word-level (for example, whether individual words are correct or not) are also common (Turchi et al., 2014; Shenoy et al., 2021).
+Recently, there has been a tendency to compare metrics to human scores derived from fine-grained error analysis as they tend to correspond better to human professional translators (Freitag et al., 2021a).
+Metrics of higher granularity (e.g., word-level) provide more explanatory value than those of lower granularity, as they provide more details on which parts of the input might be translated incorrectly (e.g., Leiter, 2021; Fomicheva et al., 2021, 2022).
+Due to their prevalence, most techniques we describe in §4 target the explainability of metrics evaluated on sentence-level.
+Quality aspect This refers to the properties a metric evaluates and measures, for example, human assigned adequacy (e.g., via direct assessment, see Graham et al.
+(2016), obtained from crowd-workers), fluency, or other aspects (such as relevance, informativeness, or correctness, mostly in other NLG fields) (e.g., Sai et al., 2022; Celikyilmaz et al., 2020).
 We use this dimension in §4, as metrics that report multiple quality aspects provide more information on specific translations, i.e., they have more explanatory value for potential users.
@@
 (2021) describe a further distinction based on the task a metric is designed to solve (or how a metric is implemented) to achieve high correlation with human judgments.
-They identify the three tasks below.444Chen and Eger (2023) argue for metrics induced from natural language inference (NLI) as a conceptual framework.
+They identify the three tasks below.333Chen and Eger (2023) argue for metrics induced from natural language inference (NLI) as a conceptual framework.
 Supervised Regression/Ranking: These models use supervised regression to train a metric to predict continuous values that directly correlate with human judgments from hypothesis and ground truth tokens.
@@
 (2020) propose a ranking setup in which they minimize a loss that becomes smaller the greater the distance between embedding representations of the two sentences become.
-Especially in the MT community, metrics trained on human scores (e.g.
-COMET; Rei et al.
-2022a) have become more and more dominant, while Belouadi and Eger (2023) argue that unsupervised metrics are more widely applicable.
-Text Generation: Some metrics use the probability that a sentence (e.g., a translation) is created by a text generation system, given an input sentence (e.g., a reference sentence), as sentence-level score.
+Especially in the MT community, metrics trained on human scores (e.g., COMET ;Rei et al.
+2022a) have become more and more dominant, while Belouadi and Eger (2023) argue that fully unsupervised metrics, i.e., metrics that were constructed without reference sentences, parallel data and human scores, are more widely applicable.
+Text Generation: Some metrics use the probability that a sentence is created by a text generation system with a paraphrasing or translation task as sentence-level score.
+For example, the probability of translating a source sentence into a hypothesis sentence with an MT model can be used as score.
 Yuan et al.
-(2021) place the metric PRISM (Thompson and Post 2020) and their own metric BARTScore in this setup.
+(2021) place the metric PRISM (Thompson and Post, 2020) and their own metric BARTScore in this category.
+In contrast, some recent metrics are based on in-context learning and directly generate scores (e.g.
+Kocmi and Federmann, 2023a; Lu et al., 2023b).
 The distinction of learning objectives is important for explainable MT evaluation, as the applicability of some explainers depends on the objective.
-MT metrics also have multiple use cases; see Figure 1 (in the following, preceding numbers correspond to this figure).
-(1) One use case is to rate a single translated text, for example, in a scenario where a human user wants to check their own human555Note, however, that humans might make different types of translation errors (Specia et al.
-2018).
+Sentence-level MT metrics also have multiple use cases; see Figure 1 (in the following, numbers written in brackets correspond to this figure).
+(1) One use case is to rate a single translated text, for example, in a scenario where a human user wants to check their own human444Note, however, that humans might make different types of translation errors (Specia et al., 2018).
 Therefore, metric performance first needs to be evaluated for this use case.
-translation or a translation they received by an MT model before using it in a downstream task (e.g.
-Murgolo et al.
-2022; Zouhar et al.
-2023).
-(2) The standard use case is to rate a large number of translations to compare MT systems (e.g.
-Freitag et al.
-2022).
-(3) Another use case is to employ metrics in the training process of MT models, e.g.
-using reinforcement learning (e.g.
-Wu et al.
-2018) or optimizing by computing derivatives with Gumbel-softmax (e.g.
-Jauregi Unanue et al.
-2021).
+translation or a translation they received by an MT model before using it in a downstream task (e.g., Murgolo et al., 2022; Zouhar et al., 2023).
+This setup can also be used to filter parallel corpora for high-quality training data of new MT models (e.g., Ramos et al., 2023; Peter et al., 2023).
+(2) A second, common use case is to rate a large number of translations of different texts by one MT system to obtain a system-level score (for example by averaging sentence-level scores) (e.g., Freitag et al., 2022).
+Then, different MT systems can be compared their system-level scores that were computed on the same corpus.
+(3) Another use case is to employ metrics in the training process of MT models, e.g., using reinforcement learning (e.g., Wu et al., 2018) or optimizing by computing derivatives with Gumbel-softmax (e.g., Jauregi Unanue et al., 2021).
 (4) The fourth use case is to use the metrics in the decoding process of machine translation systems for reranking or with minimum Bayes risk decoding Fernandes et al.
@@
 The use cases can have different audiences and explainability might play a different role to them (see 2.2).
-For example, anyone might be a user that wants to rate a translation, but usually only domain experts want to rate a translation model.
+For example, anyone might be a user that wants to rate a translation, but usually only certain user groups, such as translation experts or MT experts, want to rate a translation model.
 2.2 Explainability
@@
 Based on (1), we decide on the scope of papers we include in the taxonomy in §3.
-Explainability & interpretability Prior work has noted that the terms interpretability and explainability vary in definition and are sometimes used ambiguously throughout the literature (Barredo Arrieta et al.
-2020; Jacovi and Goldberg 2020; Vilone and Longo 2021).
-In this paper, we follow the notion of passive interpretability and active explainability (for example, Barredo Arrieta et al.
-2020).
+Explainability & interpretability Prior work has noted that the terms interpretability and explainability vary in definition and are sometimes used ambiguously throughout the literature (Barredo Arrieta et al., 2020; Jacovi and Goldberg, 2020; Vilone and Longo, 2021; Freiesleben and König, 2023).
+In this paper, we follow the notion of passive interpretability and active explainability (for example, Barredo Arrieta et al., 2020).
 In particular, we adapt the definition of Barredo Arrieta et al.
@@
 (2020) in two aspects: (1) It includes external explainers (also called post-hoc methods) and (2) it includes explanations of the metric’s output.
-We note that (2) is different from explaining a metric, as output explanations may not provide insights into how a model works.
+We note that (2) is different from explaining the metric’s decision process, as output explanations may not provide insights on how the metric works.
 Still, we include (2) in our explainability definition for two reasons.
-One is that there have been two shared tasks and other existing works on metric explainability that consider the plausibility aspect of explanations only (see the following paragraph on faithfulness and plausibility) (e.g.
-Fomicheva et al.
-2020, 2021; Zerva et al.
-2022; Xu et al.
-2023).
-Thereby, explanations these works consider might be explanations of the output that do not actually describe model internals.
-We want to accommodate these works that do not check faithfulness in our definition.
-The other reason is that we see use cases of output explanations of strong metrics, i.e., plausible-only explanations (see the following paragraph on faithfulness and plausibility).
-In contrast, a metric is interpretable if it can be passively understood by humans (for example, decision trees and k-nearest neighbors) (Barredo Arrieta et al.
-2020).
+One is that there have already been multiple shared tasks and other works on metric explainability that only evaluate how plausible explanations explain the output of a metric (e.g., Fomicheva et al., 2020, 2021; Zerva et al., 2022; Xu et al., 2023).
+We want to accommodate these works in our definition.
+The other reason is that there are use cases for such explanations of the output (specifically for the goals of making metrics more accessible and semi-automatic labeling; see §2.3).
+In contrast, a metric is interpretable if it can be passively understood by humans (for example, decision trees and k-nearest neighbors) (Barredo Arrieta et al., 2020).
 Modern MT metrics are mostly based on embeddings and black box language models, making them non-interpretable.
-The papers we categorize in this work generate explanations for MT metrics and their outputs, i.e.
-they consider explainability.
-[set/.style = text opacity = 1]
-[…]
-[…]
-[…]
-[…]
-Faithfulness & plausibility An explanation is faithful if it accurately represents the reasoning process underlying a model’s prediction(s) (Jacovi and Goldberg 2020).
-In contrast, an explanation is plausible if it convinces its audience that a model’s prediction is correct (Jacovi and Goldberg 2021), i.e., if it aligns with human judgment.
-Both properties are not necessarily dependent on each other (Jacovi and Goldberg 2020).
-For example, an explanation of an MT metric score could make sense to a human without reflecting the metrics internal details and vice-versa.
-A degree of faithfulness is indispensable to all explainability works that actually want to reason about a model’s decision processes.
-Plausibility, on the other hand, might convince explainees of decision processes that do not reflect the model’s internals correctly, or incorrect model predictions.
-As such, optimizing for plausible explanations might be unethical and lead to deceiving outcomes (Jacovi and Goldberg 2021; Jin et al.
-2023).
-However, in our taxonomy, we find that most current works on explainable MT metrics do not consider faithfulness and only evaluate plausibility (see §4), often leaving unknown whether and to which degree the model’s internal process is accurately described by explanations.
-While this is a decisive current shortcoming, we argue that explanations that may be only plausible can still benefit certain audiences and use cases of MT metrics that do not target understanding the metric: (1) We assume that current high-quality metrics already work well to judge translation quality in many settings (Freitag et al.
-2022; Zerva et al.
-2022).
-Explainers that are proven to have a high plausibility to human experts, as for example measured by Fomicheva et al.
-(2021); Zerva et al.
-(2022), will likely provide strong reasoning about the translation quality when provided with a correct score by such metrics.
-We caution that this process will have two points of failure: metric and explainer, hence the overall rate of failures will be higher than for each of them individually.
-Still, we predict that this process will work reasonably well to foster the goals of metric accessibility (see §2.3), as long as users are aware of possible failures (see future work §5), similar to the problem of hallucinations in LLMs (Ji et al.
-2023b).
-(2) Language experts should be able to discern incorrect explanations.
-For them, explanations of quality scores (produced by high quality metrics) that strive to be plausible could provide means to faster evaluating and (if necessary) correcting translations, as encompassed by the goals of metric accessibility and semi-automatic labeling (see §2.3).
-(3) LLM’s have been shown to provide better results when being prompted to explain them (Wei et al.
-2022).
-Here, more plausible explanations might lead to better scores and vice-versa (see §5).
-(4) Similar to (3), language models sometimes benefit from multitask learning (Radford et al.
-2019), so employing automatically generated explanations might assist training better metrics.
-Note that (1) and (2) correspond to the first metric use case in Figure 1.
-To summarize, plausible but not faithful explanations could still be useful either to human users or intrinsically to a metric itself in several scenarios.
-Fairness & robustness In Figure 2, we show the relationship of our definition of explainability to robustness and fairness.
-Robustness measures how well a model (here, a metric) can address execution errors, erroneous and malicious inputs and unseen data (e.g Li et al.
-2023).
+The papers we categorize in this work generate explanations for MT metrics and their outputs, i.e., they consider explainability.
+Faithfulness & plausibility An explanation is faithful if it accurately represents the reasoning process underlying a model’s prediction(s) (Jacovi and Goldberg, 2020).
+In contrast, an explanation is plausible if it convinces its audience that a model’s prediction is correct (Jacovi and Goldberg, 2021), i.e., if it aligns with human judgment.
+We illustrate the properties with the following example:
+Source: Things became complicated
+Hypothesis: Die Dinge bekamen kompliziert
+Metric(Source, Hypothesis) = 0.8
+Explanation 1: The word became is wrongly translated into bekamen instead of wurden.
+However, the meaning of the translation can easily be inferred.
+Therefore, on a scale of 0 to 1, a score of 0.8 is given.
+Explanation 2: The words Die Dinge have 8 letters. Therefore, a score of 0.8 is assigned.
+In this example, the English source sentence was (wrongly) translated into the German hypothesis sentence.
+A metric grades the translation with a score of 0.8 and two explanations (here, natural language explanations) are provided.
+Comparing these explanations, explanation 1 is a more plausible description (to translation experts) of why the score of 0.8 was assigned, as it bases the score on translation errors.
+However, which of the explanations describes the metric more faithfully cannot be inferred from the text.
+For example, internally the metric could be relying on word lengths, making explanation 2 more faithful, even though it is linguistically non-sensical.
+When explainability techniques are used for metrics, faithfulness is always desirable and often, depending on the goal of the explanations, required.
+Plausibility (to translation experts), on the other hand, is required for some goals and not important for others (see §2.3).
+As a side note, “complete faithfulness” of an explainability technique may never be established because it would require capturing the full decision process, e.g., of neural networks, in human understandable descriptions.
+For example, Jacovi and Goldberg (2020) deem it as “a unicorn which will likely never be found”.
+Fairness & robustness We also categorize methods that test the robustness and fairness of a model (here, a metric) as explainability techniques.
+This is because these tests produce explanations that will help certain audiences to understand a model’s behavior better Barredo Arrieta et al.
+(2020).
+Here, robustness measures how well a model can address execution errors, erroneous and malicious inputs and unseen data (e.g., Li et al., 2023a).
 In contrast, Li et al.
-(2023) define fairness as the goal that a model should not have biases that result in unfair treatment for certain groups of people.
-As umbrella term for these concepts, we use responsible AI Barredo Arrieta et al.
-(2020).666There are many related and partly conflicting definitions and terms in explainable AI research (e.g.
-Vilone and Longo 2021) and the given definitions are not yet firmly established in the literature.
+(2023a) define fairness as the goal that a model should not have biases that result in unfair treatment for certain groups of people.555There are many related and partly conflicting definitions and terms in explainable AI research (e.g., Vilone and Longo, 2021) and the given definitions are not yet firmly established in the literature.
 For example, Sharma et al.
@@
 The main focus of our work is to build a taxonomy of methods for explainable MT metrics and the chosen definitions allow us to reasonably relate existing work.
-Barredo Arrieta et al.
-(2020) define responsible AI as a concept that encourages the systematic adoption of several AI principles, while also setting a focus on accountability and privacy.
-Methods that test for robustness and fairness can all be grouped into the given explainability definition, as their fulfillment requires explanations that will also help certain audiences to understand a model’s behavior better Barredo Arrieta et al.
-(2020).777Barredo Arrieta et al.
-(2020) group robustness under the term confidence.
 In the MT metrics literature, this overlap in terminology is apparent for techniques that check how robust metrics are to input perturbations.
-Here, some works frame these as explainability techniques (e.g.
-Sai et al.
-2021; Karpinska et al.
-2022), while others frame them as robustness tests (e.g.
-He et al.
-2022; Chen and Eger 2023); see §4.
-In Figure 2, we only let robustness and fairness half overlap with explainability to highlight that the explanatory value increases with fine-grained explanations.
-E.g., robustness could be measured coarse-grained with a single score, such as the minimal number of adversarial perturbations required to successfully attack a metric, or fine-grained, by using predefined perturbation templates (e.g.
-Sai et al.
-2021) or attackers that will attack with perturbations of specific properties (e.g.
-Chen and Eger 2023).
+Here, some works frame these as explainability techniques (e.g., Sai et al., 2021; Karpinska et al., 2022), while others frame them as robustness tests (e.g., He et al., 2023; Chen and Eger, 2023); see §4.
+Robustness could be measured coarse-grainedly with a single score, such as the minimal number of adversarial perturbations required to successfully attack a metric, or fine-grainedly, by using predefined perturbation templates (e.g., Sai et al., 2021) or attackers that will attack with perturbations of specific properties (e.g., Chen and Eger, 2023).
 While a single score can also explain whether metric behavior is robust, susceptibility scores to various properties allow to assess metrics in greater detail.
-The same is true for fairness (Sun et al.
-2022).
-Another reason that we do not fully include robustness and fairness in the explainability definition is that explainability results from testing them, but vice-versa robustness and fairness are usually not given just by having an explainable model.
-We note that common explanation types in the evaluation of robustness and fairness are counterfactual explanations, i.e., minimal perturbed input samples that change the output of a model to demonstrate its behavior Wachter et al.
-(2017); Freiesleben (2021).
-For fairness evaluation, counterfactuals from domains that carry a risk of being biased are used Sharma et al.
-(2020); Sun et al.
-(2022).
-For robustness, existing works check whether metric scores change as expected, when presented with counterfactuals in contrast to the original hypotheses (e.g.
-Sai et al.
-2021; Chen and Eger 2023); see §4 for more details.
-Some works frame these counterfactuals as adversarial attacks (e.g.
-Sharma et al.
-2020; Chen and Eger 2023), others as diagnostic tests(ets) (e.g.
-Karpinska et al.
-2022) or challenge sets (e.g.
-Freitag et al.
-2022).
+The same is true for fairness (Sun et al., 2022).
 Explanation level
@@
 techniques can be distinguished into decision-
-understanding and model-understanding (e.g. Gehrmann et al. 2020; Holzinger et al. 2022).
+understanding and model-understanding (e.g., Gehrmann et al., 2020; Holzinger et al., 2022).
 Decision-understanding methods are described as
@@
 explaining general properties of a model that steer its decisions.
-Often, decision-understanding is also referred to as local and model-understanding as global (e.g.
-Lipton 2016; Guidotti et al.
-2018; Doshi-Velez and Kim 2017; Danilevsky et al.
-2020).
-Explanation type Further, we distinguish explanation types, i.e., the output format of the explanations (e.g.
-Wiegreffe and Marasovic 2021).
+For example, a decision-understanding method could show that a metric assigned a score of 0.5, because word xyz was translated incorrectly.
+In contrast, a model-understanding method could show that a metric will always return a score of 1, given certain conditions.
+Often, decision-understanding is also referred to as local and model-understanding as global (e.g., Lipton, 2016; Guidotti et al., 2018; Doshi-Velez and Kim, 2017; Danilevsky et al., 2020).
+Explanation type Further, we distinguish explanation types, i.e., the output format of the explanations (e.g., Wiegreffe and Marasovic, 2021).
 One example are natural language explanations that describe the model output or its internal decision process with free text.
 In the case of MT metrics, this could be a sentence like “The metric returned a score of 0.6 for the translation, as it correctly captures all details of the source sentence but has severe grammatical errors at positions […]”.
-There are many possible explanation types (e.g.
-Barredo Arrieta et al.
-2020; Wiegreffe and Pinter 2019), such as, feature importance explanations or counterfactuals.
+There are many possible explanation types (e.g., Barredo Arrieta et al., 2020; Wiegreffe and Pinter, 2019), such as, feature importance explanations or counterfactuals.
 For our taxonomy, we select those that occur in our target papers (see §3).
 The definition of each explanation type is paired with the description of methods in §4.
-Model Access Most taxonomies of explainability techniques consider the required level of model access (e.g.
-Barredo Arrieta et al.
-2020; Linardatos et al.
-2021; Madsen et al.
-2022).
+Model Access Most taxonomies of explainability techniques consider the required level of model access (e.g., Barredo Arrieta et al., 2020; Linardatos et al., 2021; Madsen et al., 2022).
 Here, techniques that require no access (that can work with any black box model) are called model-agnostic.
@@
 We do not use the dimension of interpretable models in our taxonomy (see §3) as none of the state-of-the-art metrics falls into this category.
-Model-agnostic techniques require least model access, model-specific techniques, which use the gradient, embeddings or attention weights require more access and models that are explainable by design require access to the full model to generate explanations.888Note that besides interpretable models, none of the access levels inherently leads to faithful explanations, i.e.
-explanations that correctly describe the internal processes of a model.
-Hence, this property needs to be considered separately (e.g.
-Jacovi and Goldberg 2020).
-Audiences Prior works in explainability point out that different audiences require different types of explanations (e.g.
-Barredo Arrieta et al.
-2020; Sharma et al.
-2020).
+Model-agnostic techniques require least model access.
+Further, model-specific techniques, which use the gradient, embeddings or attention weights require more access.
+Finally, models that are explainable by design require access to the full model to generate explanations.666Note that besides interpretable models, none of the access levels inherently leads to faithful explanations, i.e., explanations that correctly describe the internal processes of a model.
+Hence, this property needs to be considered separately (e.g., Jacovi and Goldberg, 2020).
+Audiences Prior works in explainability point out that different audiences require different types of explanations (e.g., Barredo Arrieta et al., 2020; Sharma et al., 2020).
 For example, non-experts might have difficulties understanding explanations given in the domain language of experts.
 Barredo Arrieta et al.
-(2020) distinguish model users, affected users, developers, regulatory entities and managers.
-Users of MT metrics are usually domain experts in MT that want to evaluate or improve MT systems, i.e.
-MT developers.
-Another group of users might be non-experts that want to use the models to check their own translations or machine generated translations.
+(2020) distinguish model users, affected users, developers, regulatory entities, and managers.
+Users of MT metrics are often experts in MT that want to evaluate or improve MT systems, i.e., MT developers.
+Another group of users are translation experts that want to compare MT systems and rate translations for usage in daily tasks (for example for post-editing).
+A third group of users might be non-experts that want to use the models to check their own translations or machine generated translations.
 Affected users might be the ones that receive translations that were graded by a metric or generated by a model that was supervised or tested with the metric.
@@
 In this section, we discuss goals that require specific focus for explainable MT evaluation.
-We orient ourselves on common goals proposed by Lipton (2016); Barredo Arrieta et al.
-(2020); Jacovi and Goldberg (2021) and identify four specific ones that match the literature we identify in §3 and use cases that might commonly occur in our field of research (see Figure 3).
+We orient ourselves on common goals proposed by Lipton (2016), Barredo Arrieta et al.
+(2020) and Jacovi and Goldberg (2021), and identify four specific ones that match the literature we identify in §3 and use cases that might commonly occur in our field of research (see Figure 2).
+Note that goal 1 and 4 address metric internals, while goal 2 and 3 mostly consider explanations of metric outputs (and with this, also the quality of the input translations).
 We follow with a short description of each of them below.
 Goal 1: Diagnose and improve metrics.
-By explaining why a metric predicted a certain score for a machine translation, where humans would assign a different score, developers might understand its weaknesses (e.g.
-Kaster et al.
-2021; Sai et al.
-2021).
+By explaining why a metric predicted a certain score for a machine translation, where humans would assign a different score, developers might understand its weaknesses (e.g., Kaster et al., 2021; Sai et al., 2021).
 This can enable architectural changes or the selection of different data sets, leading to metric improvement.
 Likewise, explaining the whole metric can unveil whether it follows desired general properties or might otherwise be led astray by carefully crafted adversarial inputs.
+This goal requires explanations to be faithful to some extent, such that developers can understand the actual faults of the explained metric.
+Plausibility on the other hand may be harmful and trick developers into believing that their metric is working, therefore, it should be evaluated separately (Jacovi and Goldberg, 2020).
 Goal 1 is especially important for the audience metric developers, as they need to understand the shortcomings to improve on them.
 Goal 2: Make metrics more expressive and accessible.
-A metric that assigns a score on a sentence level is difficult to understand for non-experts, e.g., non-native speakers.
-If a metric provides further information, such as which words it considers incorrect, it may become more accessible and thus foster trust (which may lead to more widespread adoption of the metric).
+A metric that assigns a score on a sentence-level is difficult to understand, especially for users who are not experienced with MT and MT metrics.
+If further information about a metric’s output is provided, such as which words it considers incorrect, the metric may become more accessible and thus foster trust (which may lead to more widespread adoption of the metric).
 Exemplary works in this direction are Fomicheva et al.
-(2021); Bao et al.
+(2021) and Bao et al.
 (2023).
-Accessible metrics might also help with use cases like translation error correction in the downstream tasks of language learning or human translation services (e.g.
-Zouhar et al.
-2023).
-In these latter two use cases, accessibility might not require faithfulness of the explanation (see §2.2).
+Accessible metrics might also help with use cases like translation error correction in the downstream tasks of language learning (non-experts) or human translation services (translation experts) (e.g., Zouhar et al., 2023).
+In these latter two use cases, a high average explanation plausibility (to translation experts) and a strong metric performance are important.
+This is because, in cases where (1) the metric output is correct and (2) the explanation of the output is plausible to an expert, the explanation very likely correctly describes the shortcomings and strengths of the translation (on the level of an expert).777To foster trust, it might be sufficient, but unethical, if explanations of any metric (also weak ones) are plausible to its users (not necessarily to experts) (e.g., Jacovi and Goldberg, 2020; Jin et al., 2023).
+The problem is that users could be tricked to believe into the performance of weak metrics or adapt the mindset of wrong explanations that have believable arguments.
+Hence, benchmarks of metrics and explanations in their correlation to human expert quality scores and explanations are important.
+Generally, users should be informed that hallucinations, where an explanation is untrustworthy, might occur.
+Faithfulness is not necessary in this scenario, because the quality of the explanation in describing the translation already follows from (1) and (2).
+However, faithfulness is still desirable, as the internal decision process of a strong metric might support giving correct explanations (see §2.2).
 Goal 2 is important for the audience of metric users.
-These can be either non-experts that want to understand specific metric outputs or developers of MT systems that want to understand if a metric is a good fit for their tasks (see Figure 1).
+These could, for example, be non-experts that want to understand specific metric outputs or developers of MT systems that want to understand if a metric is a good fit for their tasks (see Figure 1).
 Goal 3: Support Semi-Automatic labeling.
-Fine-grained annotations like word-level translation error labels (e.g.
-Fomicheva et al.
-2021) are especially difficult to obtain.
-Obtaining automatic explanations to aid human annotators could boost their efficiency tremendously (e.g.
-Desmond et al.
-2021).
+Fine-grained (human) annotations like word-level translation error labels (e.g., Fomicheva et al., 2021) are especially difficult to obtain.
+Obtaining automatic explanations to aid human annotators could boost their efficiency tremendously (e.g., Desmond et al., 2021).
 To our best knowledge, none of the current works has applied explainers in these settings for metrics.
-Semi-automatic labeling could also help with post-editing of translations (e.g.
-Specia and Farzindar 2010) to guide how sentences should be adapted.
-Goal 3 is important for metric developers that want to create such word-level translation error data sets and metric users that want to correct translations.
+Semi-automatic labeling could also help with post-editing of translations (e.g., Specia and Farzindar, 2010) to guide how sentences should be adapted.
+Goal 3 is important for metric developers that want to create such word-level translation error data sets and metric users who want to correct translations.
+The requirements for faithfulness and plausibility are the same as for goal 2.
 Goal 4: Checking for social biases.
-Learned metrics might be biased towards certain types of texts used during training (e.g.
-Sun et al.
-2022).
-Biases could be detected by observing explanations where sensitive attributes (for example, gender, political or racial aspects) are flagged to be important (Lipton 2016).
+Learned metrics might be biased towards certain types of texts used during training (e.g., Sun et al., 2022).
+Biases could be detected by observing explanations where sensitive attributes (for example, gender, political or racial aspects) are flagged to be important (Lipton, 2016).
 For example, explanations could show that a metric considers male names more important for a translation than female names in the same scenario.
@@
 Metric users, metric developers, regulatory entities and managers might want to prevent this from happening to conform with given regulations.
+The requirements for faithfulness and plausibility are the same as for goal 1.
 3 Literature Review & Selection
@@
 We conducted an initial literature search by selecting 11 “seed papers”.
-Then, we leveraged the API of Semantic Scholar999https://www.semanticscholar.org/ to query new papers citing the work of the seed papers and papers citing those papers (citation tree at depth 2).
+Then, we leveraged the API of Semantic Scholar888https://www.semanticscholar.org/ to query new papers citing the work of the seed papers and papers citing those papers (citation tree at depth 2).
 This amounted to 383 papers.
@@
 In later stages, we progressively added further work released after the initial search and related works that we identified from other sources.
-In total, we describe 29 papers, some of which are not shown in Figure 4 due to space constraints.
-Following §2.1, we consider higher granularity scores (e.g.
-word-level) as explanations for lower granularity metrics (e.g.
-sentence-level).
-However, without requiring faithfulness this definition allows to view any word-level metric as explainers for any sentence-level metric.
-Thereby, it would even include cases in which the word-level scores are determined completely independent from the sentence-level, i.e., cases where the explainer would not be faithful.
-Due to the high number of existing word-level approaches, for this explanation type, we only consider works that explicitly employ the word-level scores for explainability.
-Likewise, following §2.1, metrics that report multiple quality aspects such as adequacy or fluency provide more explanatory value on why an explanation is good or bad than metrics that return a single score.
-For the quality aspect explanation type, we make an exemplary selection of techniques derived from recent metrics that leverage generation probabilities (e.g.
-Thompson and Post 2020; Yuan et al.
-2021).
+In total, we describe 37 papers, some of which are not shown in Figure 3 due to space constraints, but discussed throughout the chapter.
+Following §2.1, we consider higher granularity scores (e.g., word-level) as explanations for lower granularity metrics (e.g., sentence-level), i.e., feature importance explanations.
+Due to the high number of existing word-level approaches, for this explanation type we only consider works that explicitly employ word-level scores to explain sentence-level scores.
+Likewise, following §2.1, metrics that report multiple quality aspects such as adequacy or fluency provide more explanatory value on why a translation is good or bad than metrics that return a single score.
+For the quality aspect explanation type, we make an exemplary selection of techniques derived from recent metrics that leverage generation probabilities (e.g., Thompson and Post, 2020; Yuan et al., 2021).
 We also include a few non-MT works into our taxonomy, which show how some aspects that are currently not addressed by the MT related works might be tackled.
-[…]
-[…]
-[category] (-0.100000-5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000);
-[…]
-[…]
-[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodequality
-aspects (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodefeature
-importance (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle nodefine-grained
-errors (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000);
-[category] (-0.100000-5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000);
-[…]
-[…]
-[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) rectangle nodeperturbation
-robustness (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[title] (-0.100000-4.300000/1.75, 0.100000+ 0.985000+ 0.100000) rectangle nodelinguistic
-properties (-0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[category] (0.100000- 5.100000/2, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node post-hoc (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000);
-[…]
-[title] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeblack-box (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
-[title] (0.100000+ 5.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodegradient (0.100000+ 5.100000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
-[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeattention
-weights (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
-[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodeembeddings (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
-[category] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000- 5.100000/2, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node intrinsic (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000+ 0.100000);
-[…]
-[title] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000) rectangle nodemodel specific (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.100000+ 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Thompson and Post 2020)
-(Yuan et al. 2021)
-(Kamal Eddine et al. 2023)
-(Fu et al.
-2023) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Kaster et al.
-2021) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Golovneva et al.
-2023)
-(Opitz and Frank 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Rei et al.
-2022b)
-(Rei et al.
-2023) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000) rectangle node(Treviso et al.
-2021) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Fernandes et al.
-2022b)
-(Kabir and Carpuat 2021)
-(Rubino et al. 2021)
-(Wang et al.
-2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Leiter 2021)
-(Tao et al. 2022)
-(Azadi et al.
-2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Eksi et al.
-2021) (0.100000+ 5.100000+ 0.100000+ 2.800000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Fomicheva et al.
-2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000) rectangle node(Leiter et al.
-2022b) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000) rectangle node(Lu et al.
-2022, 2023)
-(Xu et al.
-2023), … (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000+ 0.100000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+0.985000) rectangle node(Sai et al. 2022)
-(Karpinska et al. 2022)
-(He et al. 2022)
-(Sun et al. 2022)
-(Chen and Eger 2023)
-… (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000- 0.985000);
-[content] (0.100000, 0.100000+ 0.985000+ 0.100000) rectangle node(Kaster et al. 2021) (0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[content] (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000, 0.100000+ 0.985000+ 0.100000) rectangle node(Opitz and Frank 2022) (0.100000+ 5.100000+ 0.100000+ 2.800000+ 0.100000+ 2.800000+ 0.100000+ 5.100000, 0.100000+ 0.985000+ 0.100000+ 0.985000);
-[-] (0.100000- 0.100000- 5.100000/1.75 + 0.3, 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.300000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000+ 0.100000+ 0.985000) – (0.100000- 0.100000- 5.100000/1.75 - 0.100000+0.3, 0.100000+ 0.985000+ 0.100000) node[anchor=mid, below, rotate=-90, xshift=-6cm] decision-understanding model-understanding;
-[set/.style = draw]
-[…]
-Adequacy: 0.7
-Fluency: 1;
-[…]
-“Things became complicated”
-“Die Dinge bekamen kompliziert”;
-[…]
-“Things sensebecame/sense complicated”
-“Die Dinge sensebekamen/sense kompliziert”;
-[…]
-[…]
-Numeric: “3 Dinge bekamen kompliziert”lower score
-Order: “kompliziert die Dinge bekamen”lower score
-[…]
-(Example)[set, align=left, rectangle, dotted] at (-6.5, 5.8) Example Data:
-Ground Truth: “Things became complicated”
-Hypothesis: “Die Dinge bekamen kompliziert”
-Score: ;
-[…]
-[…]
 4 Taxonomy
@@
 This means, they do not use an external explainer that could be applied to other metrics.
-BARTScore (Yuan et al.
-2021) uses the autoregressive generation model BART (Lewis et al.
-2020) to predict the average word generation probability of a sentence , given a sentence :
+BARTScore (Yuan et al., 2021) uses the autoregressive generation model BART (Lewis et al., 2020) to predict the average word generation probability of a sentence , given a sentence :
 This formula sums over the log-probability of token being generated by BART with parameters given all previous tokens of and all tokens of .
@@
 For example, they associate the generation direction of A:source B:hypothesis with the properties coherence and fluency.
-This idea of using the generation probabilities in multiple directions was first introduced in PRISM (Thompson and Post 2020), but Yuan et al.
-(2021) formalized the connection of the directions to known quality aspects.
-DATScore (Kamal Eddine et al.
-2023) extends BARTScore by using a multilingual model and further generation directions (additionally using the source sentence).
-The metric GPTScore (Fu et al.
-2023) provides a suite of prompts for language models that emphasize specific aspects when requesting a translation or paraphrase.
-Then, similar to BARTScore, the generation probabilities are averaged for each word.
-For example, they use the prompt “Rewrite the following text to make it more grammatical and well-written: [reference or hypothesis] In other words,” to generate a paraphrase whose probabilities they use for a fluency score.
+This idea of using the generation probabilities in multiple directions was first introduced in PRISM (Thompson and Post, 2020), but Yuan et al.
+(2021)
+formalized the connection of the directions to known quality aspects.
+DATScore (Kamal Eddine et al., 2023) extends BARTScore by using a multilingual model and further generation directions (additionally using the source sentence).
+The metric GPTScore (Fu et al., 2023) provides a suite of prompts for language models that emphasize specific aspects when requesting a translation or paraphrase.
+Then, similar to BARTScore, the generation probabilities are averaged over all tokens.
+For example, they use the prompt “Rewrite the following text to make it more grammatical and well-written: [reference or hypothesis] In other words,”
+to generate a paraphrase whose probabilities they use for a fluency score.
 The example specifies reference/hypothesis to describe that the metric checks either the generation probability of the hypothesis given the reference, or vice-versa.
@@
 For example, Kaster et al.
-(2021); Opitz and Frank (2022) also assign quality aspect scores, but mostly use them for model understanding; see linguistic properties explanations for more details.
+(2021) and Opitz and Frank (2022) also assign quality aspect scores, but mostly use them for model understanding; see linguistic properties explanations for more details.
 In the example presented in Figure 4, we report fluency and adequacy as quality aspects.
 We assign a fluency score of 1 on a scale from 0 to 1, as the translation is fluent and an adequacy of 0.7 as the translation of “became” is incorrect.
-Goals: Quality aspect approaches could be used for metric accessibility and semi-automatic labeling.
-The usage for metric diagnosis is limited, as the approaches are often based on models that are explainable by design and might not be faithful (the models generate their own explanations).
-The current approaches cannot be used to detect social biases.
-Audiences: The quality aspects fluency and adequacy follow simple definitions, such that these metrics could be used by non-expert and expert users.
-Limitations: The generation directions might not fully match the quality aspects.
-Also, the usage of prompting leaves it to another black box setting whether the model picks up the desired properties or not.
-Lastly, the explanatory value of quality aspect approaches per se is rather small, as they only allow to reason, “the rated translation is good because of aspect x” and to guess about the model’s reasoning.
-As none of the current works101010Kaster et al.
-(2021); Opitz and Frank (2022) are considered at a later point.
-specifically targets explainability, they do not consider faithfulness.
-Plausibility is automatically considered in standard evaluation schemes with human judgments.
-Feature importance (Row 2 in Figure 4) Feature importance techniques assign scores to all features in the input of a metric (here the input tokens) that express their importance to the output of the metric.111111Other names include relevance scores or attribution scores.
-Sometimes the scores are used to highlight important words.
+Feature importance (Row 2 in Figure 3) Feature importance techniques assign scores to all features in the input of a metric (here the input tokens) that express their importance to the output of the metric.101010Other names include relevance scores or attribution scores.
 Fomicheva et al.
 (2022) note that high feature importance scores should correspond to translation errors, as humans pay most attention to errors when rating a translation.
-They build on the idea that when humans evaluate translations, they often focus on the errors that they can identify on a word- or phrase-level (Freitag et al.
-2021a).
+They build on the idea that when humans evaluate translations, they often focus on the errors that they can identify on a word- or phrase-level (Freitag et al., 2021a).
 Then, Fomicheva et al.
 (2022) evaluate how well feature importance is correlated with human word-level error annotations, i.e., the plausibility of the explanations.
-We note that, depending on the explainability method, these scores might be inverted, following the interpretation “errors are not important for a metric to achieve a high score”.121212Most gradient and attention based approaches return high scores for errors (e.g.
-Treviso et al.
-2021; Rubino et al.
-2021), for embedding-based approaches it depends on whether cosine-similarity or -distance is used (e.g.
-Leiter 2021; Tao et al.
-2022), for attribution based techniques using LIME (Ribeiro et al.
-2016) or SHAP (Lundberg and Lee 2017), low scores are returned for errors (e.g., as part of the baseline implementations in Eval4NLP21: https://github.com/eval4nlp/SharedTask2021/tree/main/baselines).
-Two shared tasks for the explainabiltiy of reference-free metrics have originated from the setting in Fomicheva et al.
-(2022) so far (Fomicheva et al.
-2021; Zerva et al.
-2022).
+We note that, depending on the explainability method, these scores might be inverted, following the interpretation “errors are not important for a metric to achieve a high score”.111111Most gradient and attention based approaches return high scores for errors (e.g., Treviso et al., 2021; Rubino et al., 2021), for embedding-based approaches it depends on whether cosine-similarity or -distance is used (e.g., Leiter, 2021; Tao et al., 2022), for attribution based techniques using LIME (Ribeiro et al., 2016) or SHAP (Lundberg and Lee, 2017), low scores are returned for errors (e.g., as part of the baseline implementations in Eval4NLP21: https://github.com/eval4nlp/SharedTask2021/tree/main/baselines).
+Two shared tasks for the explainability of reference-free metrics have originated from the setting in Fomicheva et al.
+(2022) so far (Fomicheva et al., 2021; Zerva et al., 2022).
 As the respective papers already give an overview of the participating methods, we will only briefly describe the methods and set our focus to describe the model access they require.
@@
 Approaches that work directly on black box models are explored in Fomicheva et al.
-(2022); Eksi et al.
-(2021); Treviso et al.
+(2022), Eksi et al.
+(2021) and Treviso et al.
 (2021).
-They leverage model agnostic explainers such as LIME (Ribeiro et al.
-2016) and SHAP (Lundberg and Lee 2017).
-These methods check how a metric’s sentence level score changes when the input sentence is perturbed at selected positions.
+They leverage model agnostic explainers such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017).
+These methods check how a metric’s sentence-level score changes when the input sentence is perturbed at selected positions.
 While the results correlate with the human annotations to some degree, techniques that use model access performed better in the shared tasks.
 Gradient based methods are explored by Fomicheva et al.
-(2022); Eksi et al.
-(2021); Treviso et al.
-(2021); Rei et al.
+(2022), Eksi et al.
+(2021), Treviso et al.
+(2021) and Rei et al.
 (2022a).
-These papers use methods like Integrated Gradients (Sundararajan et al.
-2017) to determine the effect each input token has on the output.
+These papers use methods like Integrated Gradients (Sundararajan et al., 2017) to determine the effect each input token has on the output.
 Attention values perform especially good for Treviso et al.
-(2021); Rei et al.
+(2021) and Rei et al.
 (2022b), leading to winning techniques for most language pairs in both shared tasks, i.e., producing the most plausible explanations.
@@
 In their final submission, they ensemble the best performing attention heads over all layers.
-Word-level contextualized embeddings are used by Leiter (2021); Tao et al.
-(2022); Azadi et al.
+Word-level contextualized embeddings are used by Leiter (2021), Tao et al.
+(2022) and Azadi et al.
 (2022).
@@
 For model-specific works, Rubino et al.
-(2021) jointly fine-tune their metrics on the word-level and sentence level using synthetic data, to directly output the computed word-level scores.131313Rubino et al.
+(2021) jointly fine-tune their metrics on the word-level and sentence-level using synthetic data, to directly output the computed word-level scores.121212Rubino et al.
 (2021) won the constrained track of Eval4NLP21, which did not allow for word-level finetuning.
-Kabir and Carpuat (2021) use LIME with Transquest (Ranasinghe et al.
-2020a) and merge the word-level importance scores with additional word-level scores obtained from a pure word-level model.
+Kabir and Carpuat (2021) use LIME with Transquest (Ranasinghe et al., 2020b) and merge the word-level importance scores with additional word-level scores obtained from a pure word-level model.
 After the shared tasks, Fernandes et al.
@@
 Rei et al.
-(2023) extend the work of Rei et al.
-(2022b) by using reference-based metrics (e.g.
-COMET) and importance scores of the references.
+(2023b) extend the work of Rei et al.
+(2022b) by using reference-based metrics and importance scores of the references.
 They construct a data set from the WMT21 MQM annotations Freitag et al.
@@
 Based on this, they employ the explanations for metric diagnosis, to test whether their metrics are more prone to ignoring certain types of critical errors.
-In their limitations, they note that the explainabiltiy techniques they use have shown to be faithful in some scenarios, but also note that they did not specifically test whether they are faithful for their architecture.
+In their limitations, they note that the explainability techniques they use have shown to be faithful in some scenarios, but also note that they did not specifically test whether they are faithful for their architecture.
 Finally, we note that many more metrics can provide feature-importance like explanations by design.
-Similar to the base metrics for Leiter (2021); Tao et al.
-(2022); Azadi et al.
+Similar to the base metrics used by Leiter (2021), Tao et al.
+(2022) and Azadi et al.
 (2022), most unsupervised matching metrics could be used to determine importance scores based on word- or embedding-similarities.
@@
 2020).
-In the example presented in Figure 5, high feature importance scores are assigned to words that were translated correctly, i.e. “Things/complicated” and lower scores indicate words that are wrong.
+In the example presented in Figure 4, high feature importance scores are assigned to words that were translated correctly, i.e., “Things/complicated” and lower scores indicate words that are wrong.
 Leiter et al.
@@
 This follows the intuition that the explanations carry additional information that can be beneficial to the sentence-level correlation of a metric with human judgments.
-Goals: The discussed feature importance approaches are tested for plausibility, so strong approaches could be used for metric accessibility/improvement and semi-automatic labeling.
-The usage for metric diagnosis is currently limited, as only samples are explained, and approaches have not been tested for faithfulness.
-The current approaches have not been used to detect social biases.
-Audiences: The relation of low/high scores and translation errors is intuitive, such that these techniques could be used by non-expert and expert users.
-Metric developers could use the methods to create new data sets or to directly incorporate them into metrics Leiter et al.
-(2022b).
-Limitations: For all approaches, it is not specifically tested whether the explanations are faithful.141414In different possible aspects (Jacovi and Goldberg 2020) Notably, Rei et al.
-(2023) consider that faithfulness is probably given in their case but note the missing tests as limitation.
-Second, certain translation errors cannot be easily captured by highlighting specific words.
-For example, the word-level scores cannot represent cases where the MT fails to explicitly express a grammatical feature that is omitted in the source language but is required to be explicit in the target language (for example, the use of articles when translating from Russian into German).
-Third, the explanation type does not provide correspondence between highlighted error words in the source and target language.
-Fine-grained errors (Row 3 in Figure 4) Some recent works detect errors on a word-level and leverage their number and type to compute a sentence-level score (Lu et al.
-2022, 2023; Bao et al.
-2023; Xu et al.
-2023).
-This follows the ideas of the MQM annotation scheme (Lommel et al.
-2014; Freitag et al.
-2021a) in which human experts evaluate translations in a similar manner.
+Fine-grained errors (Row 3 in Figure 3) Some recent works detect errors on a word-level and leverage their severity, number and type to compute a sentence-level score (Lu et al., 2023a, b; Bao et al., 2023; Xu et al., 2023; Fernandes et al., 2023; Kocmi and Federmann, 2023a).
+This follows the ideas of the MQM annotation scheme (Lommel et al., 2014; Freitag et al., 2021a) in which human translation experts evaluate translations in a similar manner.
 Freitag et al.
-(2021a) have shown that this annotation scheme improves over previous annotation methods such as direct assessments (Graham et al.
-2015).
+(2021a) have shown that this annotation scheme improves over previous annotation methods such as direct assessments (Graham et al., 2015).
 While the main goal of these works is to provide sentence-level scores, they are also explainable-by-design and the generated error-labels carry explanatory value (model-specific in Figure 3).
-In specific, BARTScore++ (Lu et al.
-2022) distinguishes explicit errors that are easy to detect and implicit errors.
+In specific, BARTScore++ (Lu et al., 2023a) distinguishes explicit errors that are easy to detect and implicit errors.
 They detect the explicit errors based on BART generation probabilities.
@@
 Lu et al.
-(2023) also combine fine-grained errors into a score.
+(2023b) also combine fine-grained errors into a score.
 They first request ChatGPT to generate explanations based on MQM annotations (erroneous words, error type and error majority).
 In a second step, they ask a model to calculate a score from these.
-They note the instabilities of ChatGPT, but find that their approach is strong on a system-level and generates sentence-level scores that are better than BERTScore.
+This is a form of chain-of-thought prompting (Wei et al., 2022), a prompting technique where a language model is instructed to first generate an explanation/reasoning before returning a result.
+Lu et al.
+(2023b) find that their approach is competitive to other state-of-the-art metrics, especially with newer versions of ChatGPT.
 Xu et al.
-(2023) fine-tune InstructScore, a LLaMA model (Touvron et al.
-2023) that generates a text containing error majority, error spans, error types (MQM) and an explanation for every error in an input sentence.
+(2023) fine-tune InstructScore, a LLaMA model (Touvron et al., 2023) that generates a text containing error majority, error spans, error types (MQM) and an explanation for every error in an input sentence.
 The generated texts are similar to Lu et al.
-(2023) but add more details on each error.
-Similarly, they construct a score from the count and severity of errors.
-They obtain synthetic training data from GPT4 (OpenAI 2023a), analyze failure modes of InstructScore during training and later optimize the metric on real machine translation outputs on data sets with annotated ground truth scores.
+(2023b) but add more details on each error.
+Similarly, Xu et al.
+(2023) construct a score from the count and severity of errors.
+They obtain synthetic training data from GPT4 (OpenAI, 2023b), analyze failure modes of InstructScore during training and later optimize the metric on data sets with real machine translation outputs and annotated ground truth scores.
 Their model shows state-of-the-art results for reference-free sentence-level evaluations.
 Further, the detailed error annotations can be used as fine-grained explanations.
-Differently to the previous approaches, Bao et al.
-(2023) fine-tune a seq2seq model to predict addition and omission errors on a word-level.
+Contrary to Instructscore, instead of fine-tuning, Fernandes et al.
+(2023) propose AutoMQM a prompting technique that leverages few-shot learning (Brown et al., 2020) and chain-of-thought prompting to generate MQM error annotations.
+They use it to prompt PALM 2 (Anil et al., 2023) and achieve competitive results for system- and sentence-level evaluation.
+In a similar way, Kocmi and Federmann (2023a) design the GEMBA-MQM prompting technique and prompt GPT4 in a few-shot and chain-of-thought setting to produce MQM annotations.
+A main distinction is that AutoMQM dynamically selects few-shot examples from pools of example translations and ratings based on the input.
+In contrast, GEMBA-MQM uses only a few, pre-selected examples and is therefore easier to apply to further language pairs.
+Both, AutoMQM and GEMBA-MQM calculate their final scores based on the detected errors with MQM heuristics (Freitag et al., 2021a).
+The WMT 2023 shared task on quality estimation (Blain et al., 2023) also includes a new sub-track where the submitted metrics should detect fine-grained errors and their severity on MQM data, to facilitate the development of more explainable metrics.
+The participants explore two main directions: (1) detection with GPT4 (Rei et al., 2023a)141414Another participating team KUNMT did not submit a paper.
+and (2) fine-tuning models for error span detection, either based on word-level supervision (Geng et al., 2023; Li et al., 2023b) or with multi-task objectives (Rei et al., 2023a).
+Notably the two winning submissions from Geng et al.
+(2023) and Rei et al.
+(2023a) are ensemble approaches, where Rei et al.
+(2023a) also use pseudo-references from Deepl and Google Translate (if their quality is better than a threshold, as measured by another metric) (Blain et al., 2023).
+Similarly, but not related to the shared task, Bao et al. (2023) fine-tune a seq2seq model to predict addition and omission errors on a word-level.
 In the example presented in Figure 4, a tag “sense” flags the wrong word sense of “became”.
 Notably, all fine-grained error approaches we describe are very recent, which is potentially caused by the adaptation of MQM in the WMT metrics shared tasks and the recent successes of LLMs.
-Goals: These fine-grained error approaches can be used for increased accessibility and semi-supervised labeling, as the methods model plausible explanations (some of them by testing the alignment of scores aggregated from explanations with human MQM scores).
-The usage for metric diagnosis is limited, as only samples are explained, and explanations have not been tested for faithfulness.
-The current approaches cannot be used to detect social biases.
-Audiences: Flagged translation errors are simple to understand, such that these techniques could be used by non-expert and expert users.
-Metric developers might use them to create new data sets to train new metrics.
-Limitations: The current approaches (besides Xu et al.
-2023) do not have the goal to provide explanations that make the sentence-level scores more descriptive.
-Instead, they only search to build sentence-level metrics from LLM
-outputs. As such, there might be open opportunities in leveraging the explanations provided by these models to explain to a human explainee. Another limitation are missing evaluations of faithfulness.
 Perturbation robustness (Row 4 in Figure 3) The methods described up to now addressed decision-understanding techniques Gehrmann et al.
@@
 This paragraph and the next describe model-understanding techniques (we describe these terms in §2.2).
-Many recent works evaluate the robustness of MT metrics (Sai et al.
-2021; Karpinska et al.
-2022; He et al.
-2022; Sun et al.
-2022; Chen and Eger 2023; Vu et al.
-2022).
+Many recent works evaluate the robustness of MT metrics (Sai et al., 2021; Karpinska et al., 2022; He et al., 2023; Sun et al., 2022; Chen and Eger, 2023; Vu et al., 2022).
 Robustness evaluation qualifies as model-understanding, as it usually infers properties that describe a model (see §2.2).
@@
 They show that perturbation checklists allow to pick metrics that are strong with respect to specific properties.
-For example, their experiments indicate MoverScore (Zhao et al.
-2019) would capture fluency better than BERTScore (Zhang et al.
-2020a) due to its ability to cope with the jumbling of words.
+For example, their experiments indicate MoverScore (Zhao et al., 2019) would capture fluency better than BERTScore (Zhang et al., 2020a) due to its ability to cope with the jumbling of words.
 The checklists allow for more fine-grained assessment of automatic evaluation metrics, which exposes their limitations and supports the design of better metrics.
@@
 (2022) use character-level adversarial attacks to evaluate the robustness of several BERTScore variants.
-The WMT21 (Freitag et al.
-2021b) and WMT22 (Freitag et al.
-2022) metrics shared tasks introduce a similar setting for robustness tests.
+The WMT21 (Freitag et al., 2021b), WMT22 (Freitag et al., 2022) and WMT23 (Freitag et al., 2023) metrics shared tasks introduce a similar setting for robustness tests.
 In WMT21, the organizers provided the participants with a challenge set composed of translations that were perturbed in a specific way, to test whether metrics reduce their score as expected when presented with respective phenomena.
-In WMT22, participants of a challenge set subtask (Alves et al.
-2022; Avramidis and Macketanz 2022; Chen et al.
-2022; Amrhein et al.
-2022) designed data sets posing challenges in a similar manner.
-In the example presented in Figure 5, we show tests similar to those performed by techniques of this category.
-The metric score should be decreased if numbers or word order used in the hypothesis change.
+In WMT22 and WMT23, participants of a challenge set subtask (Alves et al., 2022; Avramidis and Macketanz, 2022; Chen et al., 2022; Amrhein et al., 2022; Lo et al., 2023; Amrhein et al., 2023; Avramidis et al., 2023) designed data sets posing challenges in a similar manner.
+In the example presented in Figure 4, we show tests like those performed by techniques of this category.
+Here, a model-understanding explanation could be that the metric generally assigns lower scores when input sentences are subjected to numeric or word order perturbations.
 The work by Guerreiro et al.
@@
 They find that generation-based metrics show the least biases.
-Goals: The presented perturbation robustness approaches give detailed insights into metric weaknesses. Therefore, they are most suited to be used for metric diagnosis and improvement.
-Audiences: Metric developers can use these methods for diagnosis.
-They can also be used by metric users to check if a metric fits their use case.
-Regulatory entities might test if metrics fulfill their guidelines, e.g.
-exhibit no social biases.
-Lastly, affected users might gather evidence when treated unfairly.
-Limitations: A drawback of the perturbation-based approaches is that they require predefined error types.
-Therefore, errors lacking a definition in the evaluation scheme that would fool the machine translation metrics could be missed.
-These approaches are faithful to some degree, namely in the domains they have been evaluated in.
-For small perturbation sets it could be questionable if the model generally behaves as predicted.
-Therefore we would suggest evaluating each property on a large, general data sample.
-Still, it could be advisable to employ faithfulness tests, if possible, to achieve guarantees.
-Linguistic properties (Row 5 in Figure 4) Here, we summarize model-understanding techniques that directly summarize various linguistic properties a metric fulfills.
+Linguistic properties (Row 5 in Figure 3)
+Here, we summarize model-understanding techniques that directly summarize various linguistic properties a metric fulfills.
 So far, the only method that falls into this category for MT is Kaster et al.
@@
 Syntactic and morphological similarity are either captured to a smaller extent or are even negatively correlated with the metric score.
-Especially, MoverScore (Zhao et al.
-2019) and BERTScore (Zhang et al.
-2020a) have a comparatively high coefficient for the lexical score.
-As shown in Figure 5, these techniques might give explanations like “the metric puts a high focus on lexical properties (green bar) and a low focus on semantics (red bar)“.
+Especially, MoverScore (Zhao et al., 2019) and BERTScore (Zhang et al., 2020a) have a comparatively high coefficient for the lexical score.
+The example in Figure 4 shows that these techniques can give explanations that highlight which properties a metric generally focuses on.
+In the specific example, the metric puts a high focus on lexical properties (green bar) and a low focus on semantics (red bar).
 Note that the explanation of linguistic properties is close to the decision-understanding of quality aspects.
@@
 When this would be applied to MT, it would also count towards this explanation type.
-We also list Kaster et al. (2021); Opitz and Frank (2022) with quality aspect explanations, as the metrics that measure the linguistic factors can also be computed on a sentence level.
-Goals: The discussed linguistic properties approaches evaluate general metric properties and can be used for metric debugging. As they generate sentence/metric-level property scores, the techniques
-are
-not usable for word-level semi-automatic labeling.
-When used on a sentence-level, the techniques can aid the goal of making metrics more accessible.
-While not employed here, the techniques might be usable to detect social biases by introducing a set of bias detection scores from which to learn the regressors.
-Audiences: The linguistic properties can be easily explained to non-experts, however metric selection should not solely be performed based on linguistic properties
-and
-is advised to be handled by domain experts. The methods are especially useful to debug metrics, i.e. for metric developers.
-Limitations: The approach by Kaster et al. (2021) could not explain reference-free metrics well, so plausibly requires alternate explanatory variables.
-The search for regressors may be inspired by quality aspect approaches (described earlier this section) where not a global metric score is reported but several scores (such as coherence, fluency, etc.).
-These could then compose a global MT score.
-Further, some of the “property metrics” could be considered black box variables themselves and future work might replace them by more transparent factors.
-One might also explore the collinearity of the different regressors.
-For Kaster et al.
-(2021), faithfulness is likely given to some degree, around the samples that was trained on.
-Still, it could be advisable to employ faithfulness tests, if possible, to achieve guarantees.
+We also list Kaster et al. (2021) and Opitz and Frank (2022) with quality aspect explanations, as the metrics that measure the linguistic factors can also be computed on a sentence-level.
 5 Future Work
-[set/.style = draw]
-[…]
-“Die Dinge wurden kompliziert”Score:1
-“Die Dinge erhalten kompliziert”Score:0.5;
-[…]
-In German, the verb ”bekommen” means ”to get” or
-”to receive,” but in this context, it doesn’t convey the
-intended meaning effectively. Instead, we can use the
-verb ”werden” which means ”to become” to better
-capture the sense of change or transformation.;
-[…]
-“became” Source & “werden” Hypothesis
-Score;
-[…]
-Importance((“became”,“bekommen”)): 0.1
-Importance((“Things”,“Die Dinge”)): 0.6;
-[…]
-Neuron 1423: bekommen
-Neuron 5234: werden;
-[…]
-Explanation Types;
-[…]
-Ground Truth: “Things became complicated”
-Hypothesis: “Die Dinge bekamen kompliziert”
-Score: ;
-[[sep=6.2pt]-Computer Modern Rightarrow[length=3mm, width=3mm], line width=0.7mm, shorten ¡=8pt, shorten ¿=8pt] (A) edge (F) (B) edge (F) (C) edge (F) (D) edge (F) (E) edge (F) ;
 In this section, we describe future research directions for explainable MT metrics.
@@
 The other methods mostly explain in black box settings or are based on metrics that are explainable by design.
-We suppose that gradients, attention weights and embeddings have so far mostly been applied in the field of feature importance, as there is often already a 1-1 relationship their values have towards input features.
-Also, due to the two recent shared tasks, feature importance has likely received the most attention.
+We suppose that gradients, attention weights, and embeddings have, so far, mostly been applied in the field of feature importance, as there is often already a one-to-one relationship between their values and input features.
+Also, due to the recent shared tasks, feature importance has likely received the most attention.
 Future work could consider enhancing quality aspect explanations — Row 1 in Figure 3 — by incorporating more model access.
@@
 For example, analyses of the contextualized embedding space and how it affects a metric score could explain the metric on a more general level.
-Many works consider the probing of transformer-based architectures (e.g.
-Tenney et al.
-2019; Liu et al.
-2019; Sajjad et al.
-2022b) and it would be interesting to see if their results can inform the metric development.
+Many works consider the probing of transformer-based architectures (e.g., Tenney et al., 2019; Liu et al., 2019; Sajjad et al., 2022b) and it would be interesting to see if their results can inform the metric development.
 Also, current perturbation robustness techniques often use manually defined perturbation strategies.
 It would be interesting to use adversarial attacks instead that automatically infer these strategies.
-Exploring other explanation types Besides the works in our taxonomy, other fields of NLP and explainable AI in general have proposed further explanation types, e.g.
-counterfactuals, natural language explanations and rule extraction (e.g.
-Madsen et al.
-2022).
+Exploring other explanation types Besides the works in our taxonomy, other fields of NLP and explainable AI in general have proposed further explanation types, e.g., counterfactuals, natural language explanations and rule extraction (e.g., Madsen et al., 2022).
 We illustrate these options with an example in Figure 5.
@@
 (2015, 2016).
-Natural language explanations are explanations in free text form (e.g.
-Wiegreffe and Marasovic 2021).
+Natural language explanations are explanations in free text form (e.g., Wiegreffe and Marasovic, 2021).
 This is an explainability type where the generative abilities of recent large language models (LLM), like ChatGPT and GPT4, can be leveraged.
 The works by Fu et al.
-(2023); Lu et al.
-(2023); Kocmi and Federmann (2023) already explore the use of LLMs as metrics and extract the metric scores from LLM outputs (or generation probabilities) that contain (or could easily be extended to provide) natural language descriptions of why a certain score is assigned.
-In contrast, Xu et al.
-(2023) explicitly highlight the explainability aspect of their model.
-In our taxonomy, we grouped them under fine-grained error explanations, as this is their main goal.
-However, they also produce natural language explanations for each error.151515We again note that these techniques will not necessarily describe a metric’s internal workings, but might be usable due to providing plausible explanations, see §2.2.
-In these works, the explanations are rather used as a byproduct, while an evaluation of the explanation quality and further use cases remains for future work.
-The example in Figure 6 was generated with the help of ChatGPT, which shows that such an approach might yield plausible explanations.
-An example from other NLP domains is Rajani et al.
-(2019), who generate explanations for commonsense reasoning.
+(2023), Lu et al.
+(2023b), Kocmi and Federmann (2023b), Kocmi and Federmann (2023a), Fernandes et al.
+(2023) and Xu et al.
+(2023) already explore the use of LLMs as metrics and extract the metric scores from LLM outputs (or generation probabilities) that contain (or could easily be extended to provide) natural language descriptions of why a certain score is assigned.
+In our taxonomy, we grouped them under fine-grained error explanations, as this is their main goal.151515We again note that these techniques will not necessarily describe a metric’s internal workings, but might be helpful due to providing plausible explanations, see §2.2.
+In these works, the explanations are mostly used as a byproduct, while a thorough evaluation of the explanation quality and further use cases remains for future work.
+The example in Figure 5 was generated with the help of ChatGPT, which shows that such an approach might yield plausible explanations.161616We used ChatGPT in June 2023 with following prompts and asked to rephrase the output once:
+Ground Truth: “Things got complicate”
+Hypothesis: “Die Dinge bekamen kompliziert”
+Is this translation of the ground truth correct? Why or why not?
+An example from other NLP domains is Rajani et al. (2019), who generate explanations for commonsense reasoning.
 In NLG in general, many other approaches for LLM-based metrics have been constructed since the release of ChatGPT.
-These could potentially be extended to MT metrics and explainability (e.g.
-Chiang and yi Lee 2023; Wang et al.
-2023; Ji et al.
-2023a; Chen et al.
-2023; Liu et al.
-2023).
-Recent works, describe prompting strategies like chain-of-thought (Wei et al.
-2022) and tree-of-thought (Yao et al.
-2023) where language models show performance improvements if they are giving explanations of their output.
-It could be an interesting research direction, to see if the explainability techniques summarized in this paper can improve metric performance in this way, or if these techniques can provide reasonable explanations (that are at least plausible).
+These could potentially be extended to MT metrics and explainability (e.g., Chiang and Lee, 2023; Wang et al., 2023; Ji et al., 2023; Chen et al., 2023; Liu et al., 2023).
+As for example Kocmi and Federmann (2023a) point out, it it dangerous for academia to rely on proprietary models, especially for evaluation purposes.
+This includes the following issues: Metrics based on proprietary LLMs (1) could be influenced by the LLM providers, (2) could change or become unavailable at any time, (3) could carry more unexpected biases due to unknown training data and (4) could be contaminated in their evaluation by being trained on common test sets.
+Recent works, describe prompting strategies like chain-of-thought (Wei et al., 2022) and tree-of-thought (Yao et al., 2023) where language models show performance improvements if they are giving explanations of their output.
+While some of the works that are described in our taxonomy already employ such techniques to improve metrics a systematic exploration of the effect on metric scores, or whether these techniques can provide reasonable explanations (that are at least plausible), remains.
+A first work that considers this systematic exploration is the EVAL4NLP 2023 shared task (Leiter et al., 2023) where participants should improve metrics scores only by prompting.
+Their results on a small scale experiment indicate that the participants’ generated natural language explanations are not helpful to humans, in most cases.
 Neuron Explanations are another potential use case of LLMs to separately explain each Neuron in the underlying networks.
@@
 This allows to successfully simulate the behavior and effects of some neurons.
+Interestingly, they also find error related neurons, like a “post-typo-neuron”.
 For MT metrics, this would allow to determine the role certain neurons play to determine scores when specific features like “became is used” are present in the metric input.
-At the moment, these methods however still have some limitations Bills et al.
-(2023).
-Interactive Explanations Due to their context size, instruction fine-tuning and emergent abilities of recent LLMs allow to interactively ask questions and adapt based on the answers (e.g.
-OpenAI 2023b).
+At the moment, the success of neuron explanation methods is still limited (Bills et al., 2023).
+Interactive Explanations: Due to their context size, instruction fine-tuning and emergent abilities of recent LLMs allow to interactively ask questions and adapt based on the answers (e.g., OpenAI, 2023a).
 Jacovi et al.
@@
 For example, a user could ask for more detailed information for a mistranslation that was noted in a first explanation.
-New approaches could furth consider displaying multiple explanation types together, like Xu et al. (2023) who present fine-grained error annotations with natural language explanations.
+New approaches could further consider displaying multiple explanation types together, like Xu et al. (2023) who present fine-grained error annotations with natural language explanations.
 Future evaluation and concepts Current work scarcely performs human evaluation of explanations.
 This is partly because the focus is often not on the explanations but other aspects, like the sentence-level scores.
-One form of human evaluation (for plausibility) is often to use pre-computed human explanations and check the correlation to them (e.g.
-Zerva et al.
-2022).
+One form of human evaluation (for plausibility) is often to use pre-computed human explanations and check the correlation to them (e.g., Zerva et al., 2022).
 Current works on MT metrics do, however, not evaluate whether the explanations are actually helpful to explain the concepts they describe to their audience.
@@
 (2023).
-Hence, we encourage future work to address this, e.g.
-by considering simulatability Hase and Bansal (2020).
-Regarding plausibility, it could also be interesting to evaluate whether plausible explanations of incorrect metric scores can actually deceive non-experts or experts (with the current state-of-the-art), as warned by Jacovi and Goldberg (2021); Jin et al.
-(2023).
-As noted throughout the paper, the degree of faithfulness an explanation has to the metric’s internal decision process is an important property (Jacovi and Goldberg 2020), especially for metric debugging.
+Hence, we encourage future work to address this, e.g., by considering simulatability Hase and Bansal (2020).
+Human evaluation may also be performed in downstream tasks, by checking performance with and without explanations.
+Regarding plausibility, it could also be interesting to evaluate whether plausible explanations of incorrect metric scores can actually deceive non-experts or even translation experts, as warned by Jacovi and Goldberg (2021) and Jin et al.
+(2023) (for general explainability use cases).
+As noted throughout the paper, the degree of faithfulness an explanation has to the metric’s internal decision process is an important property (Jacovi and Goldberg, 2020), especially for metric debugging and bias detection.
 Current works on decision-understanding of MT metrics, however, mostly dismiss faithfulness.
 Future work might consider the evaluation of metric faithfulness, to enable these use-cases.
-We note that this evaluation will be complex, due to faithfulness being multi-faceted (Jacovi and Goldberg 2020).
+We note that this evaluation will be complex, due to faithfulness being multi-faceted (Jacovi and Goldberg, 2020).
 For example, lower-granularity scores can faithfully explain sentence-level scores that are aggregated from them, however this does not explain how the actual black box component produced the lower-granularity scores.
-Further, there are pitfalls that exist in faithfulness evaluation (Ju et al.
-2022), warning against the usage of common methods like area over the perturbation curve scores (AOPC) (Samek et al.
-2017).
+Further, there are evaluation pitfalls that exist in faithfulness evaluation (Ju et al., 2022), ruling out common methods like area under the perturbation curve scores (AOPC).
 Therefore, future work should first consider what constitutes faithful explanations in terms of metrics and design evaluation approaches.
-Further, we note that current works mostly consider explaining sentence-level scores. Future work, might also consider other metric granularity types more.
+Further, we note that current works mostly consider explaining sentence-level scores.
+Future work, might also consider other metric granularity types more, such as word-level and document-level metrics.
+Additionally, future work could explore the influence of language resource availability on the explanations (although it may largely depend on metric quality) and whether some explainable metric setups are not applicable in low-resource settings.
 Exploring domain boundaries Future work might also consider under-explored use cases of explainability for MT metrics.
@@
 (2022) and none considers the usage of explainability for semi-automatic labeling of training data.
-Besides examining just the metrics, future work could also inspect the relation of explainable MT metrics, especially plausible-only explanations, to explainable MT (e.g.
-Stahlberg et al.
-2018), as explanations of both will probably be similar.
+Besides examining just the metrics, future work could also inspect the relation of explainable MT metrics, especially plausible-only explanations, to explainable MT (e.g., Stahlberg et al., 2018), as explanations of both will probably be similar.
 Lastly, future work could explore providing similar taxonomies as ours that are comprising the whole field of explainable natural language generation metrics.
 6 Related Work
-A large number of conceptual works and surveys have been conducted in the area of explainability (e.g.
-Lipton 2016; Biran and Cotton 2017; Adadi and Berrada 2018; Doshi-Velez and Kim 2017; Došilović et al.
-2018; Guidotti et al.
-2018; Arya et al.
-2019; Carvalho et al.
-2019; Gilpin et al.
-2019; Miller 2019; Barredo Arrieta et al.
-2020; Bodria et al.
-2023; Linardatos et al.
-2021).
+A large number of conceptual works and surveys have been conducted in the area of explainability (e.g., Lipton, 2016; Biran and Cotton, 2017; Adadi and Berrada, 2018; Doshi-Velez and Kim, 2017; Došilović et al., 2018; Guidotti et al., 2018; Arya et al., 2019; Carvalho et al., 2019; Gilpin et al., 2018; Miller, 2019; Barredo Arrieta et al., 2020; Bodria et al., 2023; Linardatos et al., 2021).
 There are few surveys that touch upon the need for explainable NLG (here MT), thereby motivating our work.
@@
 Our focus on concepts for explainable MT metrics is a different one and our taxonomy of approaches is much more extensive.
-General surveys on explainable NLP are related to our work, as they highlight how explainability techniques have been used in NLP and might further guide the development of future approaches for explainable MT metrics (e.g.
-Vijayakumar 2023; Sajjad et al.
-2022a; Gurrapu et al.
-2023; Lertvittayakumjorn and Toni 2021; Saha et al.
-2022; Sun et al.
-2021; Zini and Awad 2022; Qian et al.
-2021; Madsen et al.
-2022).
-Besides these works, there are also surveys of robustness (e.g.
-Goyal et al.
-2023) and fairness (e.g.
-Blodgett et al.
-2020), which are related, as discussed in §2.2.
+General surveys on explainable NLP are related to our work, as they highlight how explainability techniques have been used in NLP and might further guide the development of future approaches for explainable MT metrics (e.g., Vijayakumar, 2023; Sajjad et al., 2022a; Gurrapu et al., 2023; Lertvittayakumjorn and Toni, 2021; Saha et al., 2022; Sun et al., 2021; Zini and Awad, 2022; Qian et al., 2021; Madsen et al., 2022).
+Besides these works, there are also surveys of robustness (e.g., Goyal et al., 2023) and fairness (e.g., Blodgett et al., 2020), which are related, as discussed in §2.2.
 The taxonomy graphic in Figure 3 was adapted from the survey of post-hoc explainability methods in NLP by Madsen et al.
@@
 A major weakness of the current realization of (1) is that error highlights do not consider the correspondence between words in source and target.
-Further, a weakness of (2) is that perturbation types have to be manually defined, an issue that might be tackled by using more general approaches of adversarial attacks in the future (we discuss this issue in more detail in an earlier version of this paper; Leiter et al.
-2022a).
+Further, a weakness of (2) is that perturbation types have to be manually defined, an issue that might be tackled by using more general approaches of adversarial attacks in the future (we discuss this issue in more detail in an earlier version of this paper; Leiter et al., 2022a).
 Generally, most current evaluations of explainability approaches for MT metrics do not consider the faithfulness of explanations, limiting their use in metric debugging and high-risk scenarios.
-We also present a vision of next-generation approaches to explainable evaluation metrics, which should help fix the problems of the above paradigms and provide guidance to explore unexplored areas.
+We also present a vision of future approaches to explainable evaluation metrics, which should help fix the problems of the above paradigms and provide guidance to explore unexplored areas.
 Here, we also urge future work to consider the desiderata and implications of faithfulness for explainable MT metrics.
@@
 In this context, we hope our work will catalyze efforts on the topic of explainable evaluation metrics for machine translation.
-Acknowledgements
+Acknowledgments and Disclosure of Funding
 Since November 2022 Christoph Leiter is financed by the BMBF project “Metrics4NLG”.
@@
 Marina Fomicheva mainly contributed to this work until April 2022.
-Wei Zhao is supported by the Klaus Tschira Foundation, Heidelberg, Germany.
+Wei Zhao was supported by the Klaus Tschira Foundation and Young Marsilius Fellowship, Heidelberg, until December 2023.
 Yang Gao mainly contributed to this work before he joined Google Research in December 2021.